### v1

# Hongyi Du1, Jiarui Yao1, Xiaocheng Yang1, Denghui Zhang1,3, Yunzhu Li4, Heng Ji1 1

Abstract
Language model agents excel in long-session planning and reasoning, but existing benchmarks primarily focus on goal-oriented tasks with explicit objectives, neglecting creative adaptation in unfamiliar environments. To address this, we introduce EscapeBench—a benchmark suite of room escape game environments designed to challenge agents with creative reasoning, unconventional tool use, and iterative problem-solving to uncover implicit goals. Our results show that current LM models, despite employing working memory and Chain-of-Thought reasoning, achieve only 15% average progress without hints, highlighting their limitations in creativity. To bridge this gap, we propose EscapeAgent, a framework designed to enhance creative reasoning through Foresight (innovative tool use) and Reflection (identifying unsolved tasks). Experiments show that EscapeAgent can execute action chains over 1,000 steps while maintaining logical coherence. It navigates and completes games with up to 40% fewer steps and hints, performs robustly across varying difficulty levels, and achieves higher action success rates with more efficient and innovative puzzle-solving strategies. All the data and codes are released [^1] https://github.com/qiancheng0/EscapeBench.
[…]
[Uncaptioned image] EscapeBench: Pushing Language Models to Think Outside the Box
Cheng Qian1, Peixuan Han1, Qinyu Luo2, Bingxiang He, Xiusi Chen1, Yuji Zhang1, Hongyi Du1, Jiarui Yao1, Xiaocheng Yang1, Denghui Zhang1,3, Yunzhu Li4, Heng Ji1 1University of Illinois Urbana-Champaign, 2Johns Hopkins University, 3Stevens Institute of Technology, 4Columbia University {chengq9, hengji}@illinois.edu
1 Introduction
Building robust language model (LM) agents for planning and reasoning has always been challenging. Recent efforts have explored how agents could compress and utilize memory (Wang et al., 2023a; Hu et al., 2023; Liu et al., 2023b; Liang et al., 2023b; Wang et al., 2024d, c; Zhong et al., 2024), perform complex reasoning (Wei et al., 2022; Kojima et al., 2022; Zhou et al., 2023a; Lin et al., 2024; Yao et al., 2023), planning (Wang et al., 2023b; Liu et al., 2023a; Hao et al., 2023; Yao et al., 2024; Zhou et al., 2024a), and reflection (Madaan et al., 2024; Zhang et al., 2024a, b; Miao et al., 2024; Dhuliawala et al., 2024) to improve task success rate. Integrating these capabilities, recent lines of work begin to build agents for embodied actions (Zheng et al., 2024; Huang et al., 2024a; Zhu et al., 2023) and tool use  (Schick et al., 2023; Qin et al., 2023; Qian et al., 2024) grounded in environments including the Web (Nakano et al., 2021; Furuta et al., 2024; Gur et al., 2024), games (Guo et al., 2023; Xu et al., 2023; Hu et al., 2024), and society (Park et al., 2023; Liu et al., 2023c; Li et al., 2023; Ren et al., 2024).
Figure 1: An agent with creative thinking should adapt its observation (e.g. hard texture of wood stick) into a novel tool-use strategy (e.g. prying objects open).
The surge of agent systems also accelerates the development of simulation environments to evaluate an agent’s capabilities, including tasks like computer-based operations (Yao et al., 2022; Deng et al., 2024; Zhou et al., 2024b; Xie et al., 2024; Liu et al., 2024b), scientific research (Wang et al., 2022; Bran et al., 2023; Boiko et al., 2023; Huang et al., 2023a), and interactive experiences in text-based (Côté et al., 2019; Urbanek et al., 2019; O’Gara, 2023; Wu et al., 2024) or virtual sandbox game environments (Lin et al., 2023; BAAI, 2023; Wang et al., 2024a). However, most existing benchmarks are goal-oriented, emphasizing models’ planning, reasoning, and error-handling abilities, while overlooking an agent’s creativity: the capacity to think innovatively and adapt its observations to new, emergent, and unstructured scenarios.
Current agents still significantly lack creativity in novel tool use (Zhang et al., 2023), as their training predominantly focuses on memorizing tool-task associations. This emphasis overshadows their ability to explore tool affordances and adapt to unstructured scenarios (Zhang et al., 2024c), both of which are critical for innovation. In robotics and autonomous systems, this limitation prevents agents from leveraging tools in unexpected ways to tackle unforeseen challenges (Bossema et al., 2023; Sandoval et al., 2022).
In fact, creativity is widely recognized as a crucial component of intelligence. In cognitive science, the well-established Triarchic Theory of Intelligence (Sternberg, 1984) divides intelligence into three components: practical, analytical, and creative. While current reasoning benchmarks primarily assess analytical intelligence through problem-solving, and simulation environments focus on practical intelligence by testing the application of knowledge in real-world scenarios, creative intelligence remains largely unaddressed. This gap is notable given its importance for enabling agents to demonstrate human-like adaptability and think beyond predefined parameters.
To bridge this gap, we introduce EscapeBench, a benchmark that evaluates LM’s creative reasoning using scenarios inspired by room escape games. These scenarios test whether a model can escape a room through trial and error by solving unconventional puzzles that require “thinking outside the box” skills including creative observation, innovative tool use, and strategic problem-solving. As shown in Figure 1, a wooden stick, typically used for walking or poking, has to be repurposed to pry open a lid due to its hard texture. This demands agents to perform adaptive reasoning under customized constraints. Overall, our benchmark sets itself apart through several distinctive features:
  • Creative Tool Use: The tools at hand might be repurposed for creative use in order to solve the puzzle. These innovative ways of tool use are uncommon in the LM agent’s existing parametric knowledge, requiring it to reason creatively and adapt its observation into customized scenarios.
  • Uncertain Goal Pathways: While the final goal of each game is escaping from the room, the pathways to achieving it cannot be explicitly foreseen. An agent cannot devise precise, long-range plans initially and must rely on trial and error to discover viable strategies.
  • Super-Long Reasoning Chain: Each scenario requires even an omniscient agent to perform over 100 steps, with at least 40 bottleneck actions required to achieve the goal. A human player may take up to an hour to complete one game.
We benchmarked multiple open- and closed-source models within the BaseAgent framework, which incorporates working memory and Chain-of-Thought reasoning (Wei et al., 2022), by objectively evaluating their performance based on the total steps and hints used to complete each game. Our results show that even the best models struggle to complete the easiest game setting without hints, often requiring up to ten times the optimal steps and falling far behind human performance. These findings highlight how models tend to be constrained by conventional thinking patterns, struggling to break free from them and show creativity.
To address this, we further introduce EscapeAgent, enhanced with Foresight for creative tool use and Reflection for implicit goal identification. The Foresight module encourages the agent to propose and self-evaluate creative hypotheses and predict outcomes for tool use before taking action, while the Reflection module helps the agent maintain an unsolved task list, providing dynamic reference to enhance its awareness of goals in future actions. Experiments show that EscapeAgent reduces hint reliance by nearly 50% and significantly lowers the number of action steps required to complete the game. It performs robustly across different difficulty levels, making progress more efficiently and exploring creative strategies with a higher success rate. In summary:
  • We identify challenges in LLM agent creative intelligence and introduce EscapeBench, a robust environment for evaluating agent creativity.
  • We present EscapeAgent, which boosts creative reasoning by identifying implicit goals and generating innovative hypotheses.
  • We propose measuring creativity through tool use and crafting, and introduce new metrics that provide a fresh dimension for agent evaluation.
2 Related Work
Creativity in Language Models.
Creativity is a cornerstone of human intelligence and a growing focus in AI research Legg and Hutter (2007); Lake et al. (2017). LMs have demonstrated notable creative capabilities across domains - they excel at generating narratives and poetry (Brown et al., 2020; Akoury et al., 2020), show effectiveness in tool creation and design (Qian et al., 2023; Cai et al., 2024), and augment human creativity through interactive ideation (Mialon et al., 2023). In scientific discovery, research has also found that LM-generated ideas tend to be more novel but slightly less feasible than those from human experts (Si et al., 2024; Wang et al., 2024b).
However, research on LM creativity still remains nascent, emphasizing novelty, surprise, and practical value through psychological assessments like the Alternative Uses Test (AUT) (Guilford, 1967) and Torrance Tests of Creative Thinking (TTCT) (Boden, 1998). Creativity in LMs is categorized as combinatorial, exploratory, or transformational, with transformational being the most challenging (Franceschelli and Musolesi, 2023). A TTCT study found GPT-4 performing in the top 1% of human fluency and originality, but adapting such assessments to other LMs faces limitations like sample randomness and high evaluation costs (Guzik et al., 2023). Similarly, a modified Torrance Test (Zhao et al., 2024) identified strengths in elaboration and originality but highlighted gaps influenced by prompts and role-play. Notably, most research evaluates backbone models, whereas our work explores creativity within an LM agent-based setting that requires complex reasoning and planning.
Agent Evaluation in Simulated Environment.
Existing work in agent evaluation primarily focuses on text-based and sandbox-like environments that assess an agent’s cognitive and behavioral abilities in goal-oriented tasks Zhou et al. (2023b); Chen et al. (2022, 2024); Yu et al. (2024); Deng et al. (2024), with emerging research exploring the integration of LM/VLM-enabled agents in robotics to address challenges in real-world, embodied settings Liang et al. (2023a); Huang et al. (2023b, 2024b); Rana et al. (2023). Text-based environments Yuan et al. (2018); Côté et al. (2019), such as interactive fiction games Lin et al. (2024) or conversational agents Qiao et al. (2023), evaluate natural language understanding, reasoning, and decision-making consistency Uludağlı and Oğuz (2023); Qi et al. (2024). Games like Zork Infocom (1980) and TextWorld measure narrative comprehension and problem-solving in structured contexts. In contrast, sandbox environments Lin et al. (2023); Gan et al. (2021); Fan et al. (2022) like Minecraft Zhu et al. (2023) and Roblox Rospigliosi (2022) provide more open-ended settings that test spatial reasoning, planning, and collaboration Carroll et al. (2019); Agashe et al. (2023). Evaluation of agents in these environments typically relies on task-specific, cognitive, behavioral, and learning metrics to assess their ability to achieve objectives, reason adaptively, and collaborate. However, existing benchmarks often neglect creative adaptation in unfamiliar environments requiring unconventional tool use and proactive problem-solving. To address this gap, we introduce EscapeBench which evaluates agents’ creative reasoning and strategic problem-solving through navigating uncertain goal-achieving pathways, offering a new dimension of agent evaluation.
Figure 2: An illustration of Scenes, Tools, and Items in the game engine design, and their relations with the model agent’s action space. Tools can be collected and used to craft new tools or applied to other items, while an item typically requires a click, input, or tool application to trigger specific effects.
3 EscapeBench Construction
3.1 Motivation
Most current agent benchmarks emphasize explicit, goal-oriented tasks, where decisions rely on commonsense knowledge encoded within the agent model’s parametric space. For instance, a typical task in AgentBench (Liu et al., 2024b) might ask: “What musical instruments did the Minnesota-born Nobel prize winner play?” Using the Freebase search API, the agent can chart a clear pathway to the goal, identifying relevant data and filtering out irrelevant results. This process is deterministic and fact-based, with each step reflecting the agent’s analytical skill in processing search outputs. While such benchmarks effectively test an agent’s information-seeking and retrieval capabilities, they often overlook its ability to engage in creative problem-solving—an area that extends beyond analytical and practical intelligence.
This leads to our core research question: How to build an environment that benchmarks an agent’s creative intelligence? To answer this, an agent must demonstrate out-of-the-box thinking in its reasoning. Since the primary function of agents lies in their tool-use abilities, we aim to design a benchmarking environment that emphasizes creative tool use. Room escape game scenarios, by nature, are well-suited for this purpose, as they require a player to creatively make use of all the things at hand to solve difficult puzzles. In the following sections, we introduce the components of our game engine (Section 3.2), the action space available to agents (Section 3.3), the data annotation process (Section 3.4), and preliminary case study results (Section 3.5).
3.2 Engine Design
Our engine aims to simulate the room escape environment that i) receives agent actions and ii) makes corresponding environment responses as agent action’s reward. It builds up a logical skeleton that ensures a smooth benchmarking process once the game configuration is given. Specifically, our game engine involves three key components:
  • Scenes: The container of tools and items, connected with each other forming a graph structure that constitutes the whole game scenario.
  • Items: Objects that are intractable in each scene. Tools, inputs, and other interactions may be applied to trigger its state change or other effects.
  • Tools: Objects that could be collected in each scene, usually applied to other items to take effect or to other tools to craft new ones.
The interaction of these three key components forms a game scenario’s basic logic. For the game engine example in Figure 2, different scenes inside a van connect with each other forming a graph to represent their physical connectivity via connectors, such as doors or tunnels. The tool key chain collected in a scene will be placed in the bag for future use, while the item wire iron is waiting for something sharp to cut it off to trigger some effects. Please refer to Appendix A for more detailed examples and explanations.
3.3 Action Space
The model agent could take five different actions in each scene. While the action space is well-defined, the parameter space—regarding the scenes, items, or tools involved in these actions—is high-dimensional, thus enabling dynamic interactions.
  • Move (Scene): Move to an adjacent scene.
  • Click (Item): Click to simply interact with an item in the scene.
  • Apply (Tool, Item): Apply a tool in the bag to an item in the scene.
  • Input (str, Item): Input an arbitrary string to an item in the scene.
  • Craft (Tool, Tool): Use two tools in the bag to craft a new one.
Figure 2 illustrates the connections between the components of the game engine and the model agent’s action space. Among all the actions, “Apply” and “Craft” stand out as the most creativity-driven, as they require the agent to think innovatively about how to use or craft tools in an unseen way during its training. We delve into specific examples in Section 3.5.
Table 1: Rules and examples for different difficulty levels. Desc. refers to Item or Tool descriptions, while Env. represents the game engine’s feedback when an unexpected action targets an Item or Tool.
Figure 3: Statistics of total Scenes, Tools, and Items across all game settings. “Key Steps” refer to the essential bottleneck actions required to complete the game.
3.4 Annotation and Statistics
Building on existing online room escape games and puzzle-solving logic [^2] https://spotlight.ee, we present EscapeBench, featuring 12 scenarios and 36 settings across three difficulty levels. Each scenario, such as a laundry room or basement exploration, is manually annotated for high-quality game engine performance. Difficulty levels are defined by the clarity of descriptions and the feedback provided for unexpected actions, detailed in Table 1. These scenarios emphasize creative tool use and crafting strategies, challenging agents throughout the game, making EscapeBench a robust environment for testing creativity. We make all our scenarios text-based to accommodate the evaluation of top-performing language models. We present in Figure 3 detailed statistics on total scenes, tools, items, and key steps involved in all game settings.
Table 2: Creative reasoning cases in EscapeBench and model’s responses versus actual creative uses of tools.
3.5 Preliminary Study
To verify the creative reasoning required in our designed games and assess the current LMs’ ability to perform such reasoning, we conduct a preliminary case study in Table 2. We sample diverse real scenarios from EscapeBench’s game design and test them on closed-source GPT-4o (Hurst et al., 2024) and open-source LLama-3.1-70B (Dubey et al., 2024). For each model, we instruct it to think creatively about what it might achieve with the provided Items and Tools in room escape scenarios.
Our results reveal that: i) EscapeBench offers a wide range of creative reasoning challenges, encompassing unconventional tool usage, implicit numerical puzzle solving, innovative tool crafting strategies, etc. ii) Both closed- and open-source models demonstrate significant limitations in creativity and imagination, particularly in identifying implicit goals and devising creative strategies. These findings emphasize the complexity and necessity of our benchmark as well as the current gaps in models’ creativity.
Figure 4: Illustration of the EscapeAgent design. Building on the BaseAgent (Action), we integrate the Foresight and Reflection modules to enhance the agent’s capabilities in creative reasoning and implicit goal identification.
4 EscapeAgent Design
Our preliminary study highlights that current LMs often struggle with identifying implicit goals and engaging in the creative reasoning needed to solve complex puzzles. To overcome these challenges and maximize the LM’s potential, we present EscapeAgent, a novel framework designed to uncover tasks dynamically and reason creatively about tool use. EscapeAgent addresses two core challenges identified by EscapeBench:
  • Uncertain Goal Pathways: We introduce a Reflection module, which dynamically maintains a task list by adding, updating, or removing tasks after each trial-and-error action. This approach fosters proactive task discovery and sharpens the agent’s focus on specific goals (Section 4.2).
  • Creative Tool Use: We design a Foresight module that facilitates explicit reasoning about creative tool applications and their alignment with specific tasks. It enables the agent to hypothesize and evaluate strategies before execution, thus promoting problem-solving (Section 4.3).
By integrating both modules with a BaseAgent, EscapeAgent excels at managing Super-long Reasoning Chains, effectively performing over 1,000 consecutive actions while maintaining logical coherence. This robust design greatly enhances the language model’s capabilities in creativity, problem-solving, and strategic thinking.
4.1 BaseAgent
The BaseAgent serves as the foundation of EscapeAgent, as shown in the middle section of Figure 4. It takes actions based on the scenario context provided by the game engine and updates its working memory with environment feedback after each step. The working memory stores the agent’s previous actions and corresponding feedback.
The BaseAgent employs Chain-of-Thought Wei et al. (2022) reasoning to decide its next action. Note that it serves as a strong baseline and standard evaluation method for EscapeBench, as it already incorporates all the essential components an agent needs for reasoning. For more implementation details, please refer to Section C.1.
4.2 Reflection Module
The Reflection Module enables the agent to maintain a structured task list, updated through:
  • New: Add a newly identified, unsolved task.
  • Update: Record attempted but failed actions.
  • Delete: Remove a task when its goal is achieved.
Each task list entry includes the task name, the target item, and previously failed actions. This organization prevents the agent from repeating ineffective actions, thus promoting efficiency.
The reflection module is triggered after every non-move action, using environment feedback to guide task list updates. For example, in Figure 4 (right), Task 1 is deleted once the machine successfully starts. By encouraging targeted problem-solving rather than random exploration, the reflection module significantly improves task-solving efficiency. See Section C.2 for more details.
4.3 Foresight Module
The Foresight Module enhances creative reasoning by explicitly evaluating tool use and problem-solving strategies. It activates in two cases:
  • When a new task is identified: The agent evaluates existing tools it has and hypothesizes potential actions to achieve the task.
  • When a new tool is collected: The agent assesses how the tool might contribute to solving existing tasks or crafting new tools.
If the module generates at least one valid action hypothesis, the agent enters a “Try Action” state, where it systematically tests the proposed actions in subsequent steps. Otherwise, it remains in a “Free Explore” state, operating like the BaseAgent to explore the environment.
For instance, in Figure 4, the foresight module identifies the action of clicking a button on the machine, thus guiding the agent into the “Try Action” state. Overall, this module enables the agent to adapt flexibly under customized scenarios, make bold hypotheses, and execute targeted action trials efficiently. See Section C.3 for further details.
5 Experiments
We divide our experiments into two parts: i) Benchmarking the creativity of current models using the BaseAgent (Section 5.2), and ii) Demonstrating the effectiveness of EscapeAgent (Section 5.3).
5.1 Settings
Environment. Experiments are conducted on 36 game settings. An agent is considered to be making progress if it either achieves a key step (bottleneck actions defined in Figure 3) or collects a new tool. Agents will receive help if they fail to make progress for 50 consecutive actions (see Section D.1), thus ensuring full completion of the game. The working memory length is set to 10 due to context window limitations.
Models. We evaluate both closed- and open-source models: GPT-4o, GPT-4o-mini (Hurst et al., 2024), Claude-3.5 (Anthropic, 2024), Gemini-1.5 (Team et al., 2024), Llama-3.1 (Dubey et al., 2024), Qwen-2.5 (Team, 2024), DeepSeek-LLM (Liu et al., 2024a), Phi-3.5 (Abdin et al., 2024), Yi (Young et al., 2024), Ministral (MistralAI, 2024). Models with fewer than 7B parameters are excluded due to near-random behavior. For consistency, we set the sampling temperature to $T=0$ and $n=1$.
Metrics. We use two main metrics including:
  • Hints Used: Total hints used in a game.
  • Total Steps: Total actions taken in a game.
Auxiliary metrics for analysis include:
  • Early Exit Progress: Proportion of key steps and tools collected before needing a hint (game progress before needing a hint for the first time).
  • Tool Hints Used (percentage): Hints used for tool collection (normalized by total tools).
  • Key Step Hints Used (percentage): Hints used for key steps (normalized by total key steps).
Results are micro-averaged across the 36 settings.
Model Name Hints Used Total Steps Early Exit  Progress (%) Tool Hints Used (percentage) Key Steps Hints Used (percentage)
GPT-4o 10.30 723.61 24.75 2.17 (8.55%) 8.14 (24.27%)
GPT-4o-mini 15.19 1002.39 16.06 2.00 (8.97%) 13.19 (38.84%)
Claude-3.5-Sonnet 8.97 690.31 28.95 1.34 (5.13%) 7.64 (22.44%)
Gemini-1.5-pro 11.06 824.31 24.18 2.50 (9.89%) 8.56 (24.83%)
Llama-3.1-70B 14.53 982.42 19.00 3.11 (12.22%) 11.42 (33.29%)
Qwen2.5-72B 16.50 1102.50 12.46 5.33 (20.97%) 11.17 (32.02%)
DeepSeek-LLM-67B 25.50 1558.47 6.63 10.50 (42.95%) 15.00 (43.73%)
Yi-1.5-34B 24.00 1573.33 11.96 8.11 (33.83%) 15.92 (46.18%)
Phi-3-medium-128k 32.19 1871.19 7.34 12.11 (49.45%) 20.11 (59.1%)
Llama-3.1-8B 25.86 1543.30 10.10 6.81 (28.56%) 19.11 (56.00%)
Ministral-8B 25.31 1556.97 8.97 7.17 (29.90%) 18.19 (53.87%)
Qwen2.5-7B 32.20 1950.42 6.52 13.81 (55.96%) 18.47 (54.43%)
Average Human 4.33 257.83 59.65 0.17 (0.69%) 4.17 (12.28%)
Oracle Chain (Upper Bound) 0.00 107.83 100.00 0.00 0.00
Table 3: Benchmarking results of baseline BaseAgent with different core models on EscapeBench. Both closed- and open-source models rely heavily on hints to complete the escape compared to human performance, with smaller-scale models exhibiting a particularly high dependency.
5.2 Benchmarking Results
We benchmark current models using the BaseAgent framework, with results in Table 3 showing that large-scale closed-source models, such as GPT-4o, Claude-3.5, and Gemini-1.5, consistently outperform smaller models. Key insights include:
  • Most hints are used on key steps, which require creative reasoning, while models can often collect tools through random exploration.
  • Models require significantly more action steps and hints than the average human player, with some using up to  20x more steps than the most efficient action chain.
Figure 5: Ablation of difficulty through Hints Used. Harder game settings generally require more hints to complete, and vice versa.
Figure 6: Ablation of difficulty through Total Steps. Harder game settings generally require more total steps to complete, and vice versa.
Complex environment causes challenges. We show in Figure 5 and Figure 6 that, generally, harder game settings require more steps and hints for an agent to solve. Since our difficulty setting depends solely on the granularity and usefulness of descriptions and feedback (see Table 1), our results demonstrate that the way the environment is presented can impact difficulty, even when the core game logic remains unchanged.
Figure 7: Distribution of Key Steps Hints Used, categorized by different actions. Colored bars represent the percentage of hints used for each action type relative to the total key steps for that type (See right of Figure 3). “Craft” and “Input” account for the highest percentage of relative hint usage, highlighting their challenge.
Input and craft are the most challenging. As shown in Figure 7, while “Apply” actions require the most hints overall, “Input” and “Craft” actions show the highest relative hint usage compared to their total occurrence in key steps. This reflects the “Input” action’s large parametric space, making random guesses impractical. Additionally, the implicit numerical challenges presented by “Input” often require synthesizing multiple pieces of evidence, adding to their difficulty. The results also suggest that the model demonstrates a degree of creativity, enabling effective “Apply” of tools, but “Craft” actions are inherently more creativity-demanding.
Table 4: Error Analysis of BaseAgent’s inefficiency or failures. All cases are selected in the same game setting. Superficial attempts dominate larger-scale models, inspiring EscapeAgent’s design to address them.
Error analysis. We observe from Table 4 that the BaseAgent often gets stuck and relies on hints due to its struggles with environment-following and creativity. Smaller models tend to perform invalid actions in complex scenarios, while larger models excel at tool collection but fail to attempt creative actions, resorting to superficial strategies. Our EscapeAgent addresses this by promoting more purposeful and creative actions.
Model Name Hints Used Total Steps Early Exit  Progress (%) Tool Hints Used (percentage) Key Steps Hints Used (percentage)
GPT-4o $5.03_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 5.27}}$ $452.75_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 270.86}}$ $47.03_{{\color[rgb]{1,0,0}\uparrow 22.28}}$ $0.33_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 1.84}}$ (1.19%) $4.70_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 3.44}}$ (13.74%)
GPT-4o-mini $10.58_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 4.61}}$ $752.25_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 250.14}}$ $28.17_{{\color[rgb]{1,0,0}\uparrow 12.11}}$ $1.14_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 0.86}}$ (4.16%) $9.44_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 3.75}}$ (28.53%)
Llama-3.1-70B-Instruct $7.92_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 6.61}}$ $645.19_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 337.23}}$ $31.44_{{\color[rgb]{1,0,0}\uparrow 12.44}}$ $1.42_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 1.69}}$ (5.44%) $6.56_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 4.86}}$ (19.15%)
Qwen2.5-72B-Instruct $9.72_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 6.78}}$ $746.61_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 355.89}}$ $28.62_{{\color[rgb]{1,0,0}\uparrow 16.16}}$ $2.72_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 2.61}}$ (10.61%) $7.00_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 4.17}}$ (20.71%)
DeepSeek-LLM-67b-Chat $20.14_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 5.36}}$ $1285.03_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 273.44}}$ $15.30_{{\color[rgb]{1,0,0}\uparrow 8.67}}$ $10.81_{{\color[rgb]{1,0,0}\uparrow 0.31}}$ (42.72%) $9.34_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 5.66}}$ (27.20%)
Yi-1.5-34B-Chat $22.59_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 1.41}}$ $1468.03_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 105.30}}$ $12.04_{{\color[rgb]{1,0,0}\uparrow 0.08}}$ $10.42_{{\color[rgb]{1,0,0}\uparrow 2.31}}$ (41.61%) $12.19_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 3.73}}$ (35.71%)
Phi-3-medium-128k-instruct $25.75_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 6.44}}$ $1513.69_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 357.50}}$ $9.99_{{\color[rgb]{1,0,0}\uparrow 2.65}}$ $11.20_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 0.91}}$ (44.99%) $14.55_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 5.56}}$ (43.35%)
Llama-3.1-8B-Instruct $19.81_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 6.05}}$ $1271.53_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 271.77}}$ $19.22_{{\color[rgb]{1,0,0}\uparrow 9.12}}$ $6.64_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 0.17}}$ (25.23%) $13.22_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 5.89}}$ (39.34%)
Ministral-8B-Instruct $19.47_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 5.84}}$ $1233.72_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 323.25}}$ $19.34_{{\color[rgb]{1,0,0}\uparrow 10.37}}$ $6.61_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 0.56}}$ (25.56%) $12.86_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 5.33}}$ (38.93%)
Qwen2.5-7B-Instruct $27.53_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 4.67}}$ $1639.58_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 310.84}}$ $8.56_{{\color[rgb]{1,0,0}\uparrow 2.04}}$ $13.00_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 0.81}}$ (52.66%) $14.58_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 3.89}}$ (43.99%)
Table 5: Benchmarking Results of our approach EscapeAgent with different core models. Nearly all the performance raises compared to Table 3, showcasing the effectiveness of EscapAgent in promoting the agent’s creativity.
5.3 EscapeAgent Results
The introduction of the Reflection and Foresight modules in EscapeAgent, as shown in Table 5, leads to a marked reduction in hint usage and total steps, with large-scale models benefiting the most. Key insights include:
  • Larger models like GPT-4o consistently outperform smaller models, suggesting that while the new modules aid creative reasoning, the core model’s capabilities remain a crucial factor.
  • Early exit progress improves significantly across models, despite the exponentially increasing difficulty of maintaining consecutive meaningful progress without hints. Notably, GPT-4o achieves near-human-level performance, demonstrating the effectiveness of the enhanced framework.
Figure 8: Distribution of step intervals for progress made through Tool Collection and Key Step achievement. EscapeAgent uses fewer steps to achieve the next progress and relies less on hints.
Figure 9: The accumulated number of completed games (in total 36) relative to total steps a game setting takes. EscapeAgent, shown in dotted lines, completes games in fewer steps, demonstrating greater efficiency.
EscapeAgent progresses more efficiently. Figure 9 demonstrates EscapeAgent’s more efficient progress, as the slopes of the dotted lines are consistently steeper. We further show in Figure 8 that EscapeAgent requires fewer actions to achieve the next valid Key Steps compared to BaseAgent, indicating better creative reasoning. The spike at 50 steps corresponds to explicit hints provided after the model’s prolonged inactivity for consecutively 50 steps, and the lower pink bar here further highlights EscapeAgent’s reduced hint dependency. Interestingly, for all models, sudden progress after 15 steps without hints is rare, highlighting a lack of spontaneous insights that humans often demonstrate, even in the EscapeAgent framework.
Figure 10: A comparison of action trial times distribution for a specific item before success. EscapeAgent tries more actions with a higher overall success rate.
EscapeAgent tries more with higher overall success. In Figure 10, we analyze the valid actions agents attempt on different items in scenes before successfully operating them. While the BaseAgent exhibits a higher first-attempt success rate, EscapeAgent achieves greater effectiveness by making multiple attempts, leading to a significantly higher overall success rate within 10 trials. This difference can be attributed to EscapeAgent’s strategy of proposing multiple viable actions simultaneously. Although this increases the likelihood of eventually succeeding by trying more actions, it does not prioritize the most-likely-to-succeed action first. As a result, the BaseAgent appears more efficient on its initial trial, but its superficial approach, as highlighted in Table 4, limits its overall performance. The EscapeAgent’s design effectively addresses this limitation by leveraging a more exploratory approach, which proves advantageous in complex scenarios requiring creativity.
Figure 11: Case study on Human, BaseAgent, and EscapeAgent’s progress map corresponding to six game settings. Although EscapeAgent uses 40% fewer hints and makes significant progress independently, it still falls far short of average human performance, often requiring twice as many total steps to complete a game.
Case study. Figure 11 presents a case study of agent progress relative to action steps across six game settings. Red dots indicate steps where hints were provided. We observe that i) EscapeAgent requires significantly fewer hints and achieves a steeper progress curve; ii) EscapeAgent demonstrates the ability to make consecutive progress in shorter intervals (e.g., the light blue and pink lines for both GPT-4o and Llama-3.1-70B); iii) harder scenarios remain challenging, especially for BaseAgent, which often relies heavily on hints to make progress (e.g., the later part of the yellow line for GPT-4o); and iv) Average human still performs far better than all agents. Humans rarely make mistakes shown in Table 4, while both agents still struggle with short memory due to context length and creative tool use strategies. These findings emphasize the need for further improvements and highlight EscapeBench’s challenge to even the most advanced models.
6 Discussions and Future Directions
LM’s creativity for benchmarking. Our experiments with EscapeBench reveal key insights into the capabilities and limitations of current language models in creative reasoning and tool use. Even the most advanced model, EscapAgent, requires more hints and takes twice as many steps as the average human to solve tasks. The benchmark demonstrates that these advanced models struggle with implicit goal identification and creative problem-solving, often relying heavily on extensive hints. This underscores a significant gap in the evaluation of AI agents: while analytical and practical intelligence are well-assessed, creative intelligence remains largely underexplored. Addressing this gap may require fundamentally improving how LMs organize knowledge, linking objects not just by words but by their affordances—how their properties and functions relate to potential uses.
Human-AI Collaboration. Human-AI collaboration in escape room puzzles holds the potential to significantly enhance problem-solving efficiency and introduce a new paradigm of learning. Humans contribute intuitive insights and creative ideas that language models are unlikely to generate independently, while AI excels in systematic tasks such as data analysis and information aggregation. Furthermore, this collaboration leverages the strength of chain-of-thought reasoning, a process through which AI organizes and evaluates information in a logical, step-by-step manner. Integrating human creativity with AI’s structured problem-solving capabilities together in the future might not only foster innovative strategies but also enable deeper learning experiences, making the problem-solving process more dynamic and efficient.
Theoretical Foundations for AI Creativity. Understanding the cognitive mechanisms behind human creativity is essential for designing AI systems that emulate or surpass human creative processes. Human creativity is a multifaceted phenomenon involving the generation of novel and valuable ideas, problem-solving, and adaptation to complex situations Dainys (2024). Neurologically, it arises from the interplay between stochastic neuronal noise and structured, learned information, driven by spontaneous brain activity fluctuations Malach (2024). This contrasts with AI’s reliance on data patterns and algorithmic processes. Boden identifies three AI creativity mechanisms: combining familiar ideas in novel ways, exploring conceptual spaces, and enabling transformative innovations Boden (1998). Our research shows that even advanced models like GPT-4o struggle with implicit goal identification and creative problem-solving, often requiring extensive prompting. The EscapeAgent framework significantly reduces this dependency and improves task-solving efficiency, indicating that these modules effectively overcome barriers to creative reasoning. However, the results also highlight the importance of the core model’s capabilities, as larger models like GPT-4o benefit more from the framework than smaller models. This suggests further research is needed to address current models’ limitations in generating truly novel ideas. Interdisciplinary approaches integrating psychology, neuroscience, and model architecture can advance agent creativity further.
Multimodal Integration. Expanding the escape room environment to include multimodal data, such as visual and voice cues, offers a promising avenue for enhancing both the agent’s performance and the realism of the scenarios. While integrating vision-language models could enable agents to interpret visual clues more naturally, such an extension would also require robust visual understanding and reasoning capabilities to handle the complexity of tasks effectively. Additionally, incorporating multimodal interactions presents opportunities to study how agents synthesize information across modalities, such as correlating visual patterns with instructions or adapting strategies based on dynamic, multimodal feedback. Future work should explore how to seamlessly integrate and harmonize these diverse data types into the benchmark, pushing the boundaries of agents’ ability to process, reason about, and act on streams of information.
Step RL for Creative Reasoning. Introducing reinforcement learning into the EscapeBench task is expected to enhance the accuracy and efficiency of the agent’s exploration in long-chain tasks. This implies that under the guidance of rewards and penalties, the agent can explore in the correct direction more quickly. Compared with existing end-to-end reinforcement learning schemes, which rely on the final completion of the escape room task as the ultimate reward, introducing step rewards—providing immediate feedback for each step of the model’s operation—could potentially accelerate convergence and foster creative advancements in the model. Specifically, the discrete steps in EscapeBench can be organized either as a continuous, logically coherent progression or as strategic, abrupt jumps reflecting non-linear reasoning. Step Rewards can also further draw on the idea of task decomposition to be structured hierarchically. By constructing such a framework, Hierarchical Reinforcement Learning could manage super-long reasoning chains, decompose complex tasks into manageable subtasks, and enable a more fine-grained exploration and evaluation of AI’s creative capabilities.
7 Conclusion
In this work, we introduce EscapeBench, the first benchmark for advancing LM’s creativity. Our results highlight that current LMs still lag in creative reasoning. However, the EscapeAgent framework effectively lowers the barrier to creative problem-solving and implicit goal identification, resulting in significant performance improvements. While these advancements are promising, challenges remain in further enhancing the creative capabilities of the language model itself. Future work may focus on integrating multi-modal perception and developing new reinforcement learning algorithms to inspire even greater creativity. Our work serves as an important first step, offering a robust environment for experimentation. Looking ahead, we believe that creative intelligence, beyond just analytical and practical capabilities, will play a key role in shaping the frontier of AI.
Limitations
Our work utilizes a text-based environment to evaluate common language models, focusing on creative reasoning within this framework. However, an Escape Room scenario inherently includes visual and auditory clues, which we have not incorporated into this benchmark. Expanding to include multi-modal inputs could be a valuable next step for future work. Additionally, while the data used in our benchmark is annotated through intensive human effort to ensure high quality, this approach limits scalability. We have explored the use of GPT-4 for automatic annotation through free exploration but found that the model sometimes overlooks important items and clues, and struggles to design environment feedback crucial for adjusting the game’s difficulty. We anticipate that more powerful vision-language models may enable better automatic annotation in the future, though current model capabilities are still a limiting factor.
Appendix
Appendix A Engine Design Details
The game engine involves scenes, tools, and items as three main components. We will introduce in detail each of them in the following.
A scene typically includes a description, its connections to other scenes, and the tools and items it contains. A typical scene example in the game configuration looks like the following:
  Scene Example
- name: hallway
desc: You are in a hallway with a blocked path straight ahead, a locked cabinet on the left, and a corridor to the right.
scene_relations:
To the blocked path close-up: blocked path close-up
To the cabinet close-up: cabinet close-up
items:
tools:
In this example, the name of this scene is “hallway”. It leads to nearby scenes including “blocked path close-up” and “cabinet close-up”, where the model could reach through action “move(To the blocked path close-up)” and “move(To the cabinet close-up)”.
Each tool has various states and a visibility status. In each state, a tool is either awaiting the application of another tool or ready to be applied to another tool or item. A typical tool example in the game configuration looks like the following:
  Tool Example
- name: key
visible: False
states:
- desc: A rusty silver key
wait_for:
- lubricant
- desc: A silver key shining bright light, ready to use now
apply_to:
- safe
In this example, a rusty key (tool, state 1) hidden in the chest (item) won’t be visible to the user until the chest is opened, and after applying lubricant (tool), it may change to a non-rusty functional key (tool, state 2) that could be applied to open a safe.
Item is an upgraded version of the tool, as each state may await multiple inputs or tools in order to trigger certain changes, including item and tool’s visibility, state, etc. A typical item example in the game configuration looks like the following:
  Item Example
- name: digital lock
states:
- desc: A digital lock linked to a card reader, power on now.
transitions:
- wait_for:
- apply, card
trigger:
- change_state, 1
reward: Authorization succeeds, you have to input a 4-digit password.
- desc: A digital lock already authorized. The password panel awaits a 4-digit input.
transitions:
- wait_for:
- input, 1672
trigger:
- change_state, item, cabinet door, 2
- change_state, 2
reward: The password is correct. A door opens somewhere ...
- desc: A digital lock. You have already input the correct password.
For instance, a digital lock (item) may await the application of card (tool) for authorization and correct password input to trigger the closed cabinet door (item, state 1) to open (item, state 2).
Appendix B Data Annotation Details
We recruited eight human annotators, all with prior Room Escape game experience (offline and online) and at least a bachelor’s degree. To ensure a smooth annotation process, all annotators were U.S.-based students with computer science backgrounds. Each annotator received detailed guidelines to ensure objective annotations and was tasked with extracting game logic and object descriptions (scenes, items, and tools) based on the official guide. The foundational data logic will be released with the software, and all annotators consented to the data collection.
Appendix C Agent Design Details
For both BaseAgent and EscapeAgent, we apply the same system prompt for its action-taking to ensure fairness:
  System Instruction
You are in a Room Escape game. You should explore the scene and find out what to do next.
There are three types of interactives: items, which are the interactable things in the scene; tools, which are applicable tools in your bag; scenes, whcih are interactable scenes near your current position.
You can perform one of the following actions:
- click(<interactable item>): Click an <interactable item> to examine it or interact with it. For example, you can examine a door handle that is marked as interactable.
- apply(<applicable tool>, <interactable item>): Apply an <applicable tool> in your bag to an <interactable item>. For example, you can apply a key in your bag to an interactable locked door to open it.
- input(string, <interactable item>): Input a string (only digits and letters) to an <interactable item>. For example, you can input a string password to an interactable combination lock.
- move(<interactable scene>): Move to a nearby <interactable item> to further explore. For example, you can move to the living room to explore more interactable items there.
- craft(<applicable tool>, <applicable tool>): Use one <applicable tool> in bag to another <applicable tool> in bag to craft something new. For example, you can use a battery in bag to a controller in bag to craft a new charged controller.
For instance, some valid actions may be: click(microwave), apply(key, silver chest), craft(controller, battery), input(c79a1, combination lock), move(Go to operation room).
The system prompts explicit instructs on the agent’s action space with examples. In the following, we present in this section more details on EscapeAgent design, including BaseAgent, Reflection, and Foresight modules.
C.1 BaseAgent Details
At each step, the BaseAgent receives information from the game engine. This information typically includes an environment description, a list of interactable objects in the scene, and the tools available in the agent’s bag. A typical environment description appears as follows:
Scene Description:
You are in the scene underneath part of the van’. There is a stepladder on the right side. There is a license plate on the left side.
Here are the items you can see in this scene:
- On the left side, there is license plate space: The license plate is currently fixed to the van, with four screws on each corner
- On the right side, there is stepladder: The stepladder is unfolded, now you can reach the top of the van,
Possible Actions:
Here are all the items in the scene that you can perform click’, apply or input’:
<interactable item> license plate space
Here are nearby scenes that you can perform move to further explore:
<interactable scene> Back to the back of the van: It leads to back of the van
Tools in Bag:
Here are the tools in your bag. You can perform craft to use two tools in your bag to craft a new one, or perfom apply to apply one tool in your bag to an object in the scene:
<applicable tool> bunch of keys: a bunch of keys with a keychain and some rust on it
<applicable tool> rag: a rag soaked with engine oil"
The scene typically includes a general description, while each item within the scene is accompanied by a detailed description. Possible actions specify which items or aspects of the scene the agent can interact with, and tools in the bag indicate which tools are available for use.
This environment description will also be coupled with working memory of previous steps. Each step’s memory contains the following fields:
History: [Step ...]
Your position: <How you get to that position from the beginning scene> e.g. Living Room -> Outside Corridor
Your action: <The action taken> e.g. move(Explore the blocked path)
Response from the environment: <Feedback from game engine> e.g. Action executed successfully. Change to another scene: blocked path close-up.
Using this information, the BaseAgent is instructed to explicitly apply a Chain-of-Thought reasoning process to determine its next action. This action is then parsed and sent back to the game engine. The game engine updates its state based on the agent’s input and provides feedback to the agent. By default, this feedback indicates whether the action was successful or not. In Easy and Normal game settings, additional customized environment feedback is provided. However, in the Hard game setting, no extra information is given.
C.2 Reflection Module Details
The Reflection module is integrated as a downstream component after BaseAgent within the EscapeAgent design. This module is responsible for maintaining a task list that is updated solely based on the agent’s current actions and the feedback received. Each task in the list generally includes the following fields:
[Task Index <index>] Name: <brief task name>, Target Item: <item name>
- Task description: <description of the task>
Example Task:
[Task Index 1] Name: open the chest, Target Item: chest
- Task description: To open the chest wth a matched key, I have tried simple click, apply safe key but all failed.
The task index facilitates the identification of specific tasks during task list management operations. The target item specifies the item in the scene that the task is focused on, enhancing the agent’s sense of purpose when exploring and performing trials within the scene. The task description outlines a potential strategy for solving the task, including actions the agent has previously attempted but failed. The following system prompt is used to guide the model:
  System Instruction
You are a helpful agent to reflect on your action and environment response, and then maintain a task list with solving suggestions.
The role of this task list is that there are some tasks you currently cannot solve with the tools at hand, but you think you may need to solve them later, so write them down with some suggestions and hints for your future reference.
After analyzing your current action and the response from the environment, you should give an action to maintain the task list: update ,new, delete or none.
- update(updated_feedback): The parameter should an updated feedback about what you newly tried but failed. The updated feedback should retain the original feedback, and add one new hindsight you got from current action.
- new(task_name, feedback): The first parameter should be a brief name of the new task you propose, the second parameter should be what you have to do (extract hint from environment response) to solve this task.
- delete(index): If you choose delete, then the first parameter should be the index of the task in the task list that you thought you have completed or is not useful anymore.
- none(): If you choose none, do not give any parameter, it indicates you believe you dont need to perform any action on the task list in the current step
For instance, valid task list maintaining action may be: update(The door has a keyhole and needs a key. I try apply a hammer but fails.), new(open the safe, I need a 4-digit password input to open it with a hint of sigma sign beside the safe.), delete(1), none().
Note that the operations “update” and “delete” can be implemented in a rule-based manner. An update is triggered when the model attempts an action on a specific item and fails, requiring revised feedback. A delete occurs when the model successfully performs an action that advances progress, necessitating the removal of the corresponding task, if it exists, from the task list.
C.3 Foresight Module Details
The Foresight module serves as an upstream component preceding the BaseAgent in the EscapeAgent design. This module is activated when a new tool is collected during the last action step or a new task is added during the previous reflection step. When a new tool is collected, the agent is provided with the current task list and instructed to propose potential applications for the tool within the context of these tasks and their specific scenarios. Additionally, the agent is given a list of all existing tools in its bag and encouraged to creatively evaluate whether the new tool could be combined with others to craft something useful. The following system prompt is employed to guide the model:
  System Instruction
You are in a Room Escape game. You have to use your creativity to figure out the use of the tool you have just collected.
There are generally two ways about how to use the tool:
1. Combine this tool with another one in your bag to craft a new tool. In this case, use acton craft(<collected tool>, <applicable tool>)’, e.g. craft(controller, battery) indicates use a battery in your bag you already have to the controller you just collected to craft a charged controller.
2. Apply this tool to a target item in a task to try solve this task. In this case, use action apply(<collected tool>, Target Item in a task)’, e.g. apply(key, locked cabinet) indicates apply the key you just collected to a locked cabinet to open it.
Here are some general hints that you may follow:
1. Please especially pay attention to the description of the task and the tool, try to find the connection between them to justify your action.
2. In your ’- Thought: ...’ part in response, you shuold explicitly think about whether theres item in bag for crafting, or task in the list for applying this tool. You should read and infer carefully from the tool descriptions and the task description, and evaluate one by one.
3. In your ’- Actions: ...’ part in response, you should give zero to multiple action calls. For each action, you should follow the format craft(<collected tool>, <applicable tool>)’ or apply(<collected tool>, Target Item in a task)’. If its a craft action, you should justify why crafting here makes sense. If its an apply action, you should first give the task index corresponding to the target item, then justify why this tool may solve the task.
If a new task is created, the agent is provided with a list of all the tools currently in its bag. It is then instructed to reason creatively about which tools could be applied to address the newly created task. The following system prompt is used to guide the model:
  System Instruction
You are in a Room Escape game. You have to use your creativity to figure out if you could use any tools you have now to solve a new task you have just discovered.
There are generally three ways to solve a task:
1. Click the target item to simply interact with it to solve the task. In this case, use action click(Target Item in current task)’, e.g. click(microwave) indicates click the microwave to examine it and try solve the task.
2. Use the tool in your bag to apply to the target item in the task. In this case, use action apply(<applicable tool>, Target Item in current task)’, e.g. apply(key, locked cabinet) indicates apply the key in your bag to a locked cabinet to open it.
3. Input a string to the target item in the task. In this case, use action input(<any string>, Target Item in current task)’, e.g. input(2413, combination lock) indicates input a string password to the combination lock to solve the task.
Here are some general hints that you may follow:
1. Please especially pay attention to the description of the task about what might be needed. Please always first try simple click to interact if havent done so. Examine the tool description and your memory pad, try to find the connection between them and what this task needs to justify your action.
2. In your ’- Thought: ...’ part in response, you should explicitly think about whether theres item to click, tool in bag for applying, or hint from memory pad and tools for string input. You should read and infer carefully from the task description, evaluate one by one.
3. In your ’- Actions: ...’ part in response, you should give zero to multiple action calls. For each action, you should follow the format click(Target Item in current task)’, apply(<applicable tool>, Target Item in current task)’, or input(<any string>, Target Item in current task)’. You shuold justify why this action may solve the task according to the task description, tool description, and memory pad hint.
Depending on whether the model proposes a valid action, the agent transitions into either the “Free Explore” or “Try Action” state. If multiple actions are proposed, the agent attempts them sequentially until a successful action is achieved. In cases where the task cannot be solved with the currently available tools, the task remains on the task list, and the newly acquired tool stays in the bag. For consistency, all action trials are included in the total count of action steps.
Figure 12: More model’s analysis on progress-making interval, extension of Figure 8.
Figure 13: More model’s analysis on item trial times, extension of Figure 10.
Appendix D More Experiment Details
D.1 Help Setting
We provide help to the agent through explicit instructions, focusing on two aspects: i) identifying the next action location that could help the model make progress, and ii) specifying the action the model should take. These objectives are addressed by providing the instruction below to the model during the action-taking step:   Action Instruction
Since youre stuck, the system will provide you with a hint. You MUST follow the hint to complete next key step.
The next target location should be: <target position>.
Your next target action should be: <target action>.
You should go to the target position and perform the target action. If you are already at the target location, please directly perform the action.
This help will remain available until the model successfully performs the target action. All the helps on how to complete the game is provided through human annotation.
In each game, there may be cases where multiple actions can be taken in parallel since they do not interfere with one another. As a result, there is no fixed sequence among them. However, to simplify the help provided, the actions are linearized into a single action chain that ensures the agent can complete the game. Whenever the agent requires help, we always provide the first action in this chain that the model has not completed, even if the model may have already succeeded in performing some later actions in the chain.
Figure 14: An illustration of progress-making trend through all 36 game settings.
D.2 Resource Setting
For closed-source models, we utilize standard APIs for testing. Under the BaseAgent framework, the average number of API calls per game setting is approximately 800, resulting in a total cost of \$50–60 per model test. While the EscapeAgent framework introduces two additional modules, these are not always triggered. Consequently, the total number of API calls increases to roughly 1.2 times that of BaseAgent, raising the cost to approximately \$60–80 per model test.
For open-source models, all benchmarks are conducted using the vLLM framework on 2 A100-80G GPUs. Inference time varies based on model size: smaller-scale models complete all 36 game settings in approximately 12 hours, while larger 70B-scale models require about twice as much time for benchmarking. The average tool-calling frequency, reflected in the Total Steps metric, is reported in Table 3 for BaseAgent and Table 5 for EscapeAgent. These metrics vary significantly depending on the specific open-source models being tested.
D.3 More Study Results
In Figure 12, we present additional results on the action step intervals with respect to two types of progress-making ways, achieving Key Step or Tool Collection. It can be observed that EscapeAgent consistently requires fewer steps to perform the next bottleneck Key Step. In contrast, for Tool Collection, the difference between the two bars is less pronounced. Despite this, the findings still demonstrate the effectiveness of our design. Tool Collection typically occurs after the successful application or crafting of tools, meaning the reasoning challenge associated with it is significantly lower. Since EscapeAgent focuses primarily on creative reasoning, it is reasonable that it excels in identifying the next Key Action more efficiently.
In Figure 13, we present the results of action trial counts for a specific item across four additional models. We observe that EscapeAgent achieves a higher success rate within 10 trials, even though it does not always succeed on the first attempt. Furthermore, for smaller models like Phi-3 and Ministral, EscapeAgent occasionally outperforms BaseAgent even in terms of one-trial success rates. This highlights how our framework effectively lowers the barriers for creative reasoning, even for less capable language models.
In Figure 15, we showcase eight additional pairs of progress-making maps for eight more models. These case studies illustrate two key points: i) There are significant disparities in creativity among models. For instance, models like GPT-4o-mini and Qwen-2.5-72B require only two-thirds of the total steps than others to achieve success, while smaller models such as Qwen-2.5-7B heavily rely on hints to make progress, even with the EscapeAgent framework. ii) When combining these results with Table 5, we observe that EscapeAgent’s performance improvement relative to BaseAgent is more pronounced for larger-scale models like GPT-4o and 70B-scale models. Conversely, while smaller models also benefit from reduced steps and hint usage, they still lag significantly behind in creative intelligence. This underscores the importance of enhancing a model’s intrinsic reasoning abilities, as our method primarily mitigates the barriers to creative reasoning but does not fully address inherent limitations.
Lastly, Figure 14 depicts the progress curves across all 36 game settings for GPT-4o and LLama-3.1-70B-Instruct. The trends reveal that BaseAgent’s total step distribution spans a broader range compared to EscapeAgent, resulting in a relatively milder and less steep progression curve. These findings further confirm the effectiveness of EscapeAgent in facilitating more efficient progress.
Figure 15: More model’s analysis on progress with respect to action steps, extension of Figure 11.

### v2

@@
-Hongyi Du1, Jiarui Yao1, Xiaocheng Yang1, Denghui Zhang1,3, Yunzhu Li4, Heng Ji1 1
+Language Model Agents
 Abstract
@@
 Experiments show that EscapeAgent can execute action chains over 1,000 steps while maintaining logical coherence.
-It navigates and completes games with up to 40% fewer steps and hints, performs robustly across varying difficulty levels, and achieves higher action success rates with more efficient and innovative puzzle-solving strategies.
+It navigates and completes games with up to 40% fewer steps and hints, performs robustly across difficulty levels, and achieves higher action success rates with more efficient and innovative puzzle-solving strategies.
 All the data and codes are released [^1] https://github.com/qiancheng0/EscapeBench.
-[…]
-[…]
-[Uncaptioned image] EscapeBench: Pushing Language Models to Think Outside the Box
+[Uncaptioned image] EscapeBench: Towards Advancing Creative Intelligence of
+Language Model Agents
 Cheng Qian1, Peixuan Han1, Qinyu Luo2, Bingxiang He, Xiusi Chen1, Yuji Zhang1, Hongyi Du1, Jiarui Yao1, Xiaocheng Yang1, Denghui Zhang1,3, Yunzhu Li4, Heng Ji1 1University of Illinois Urbana-Champaign, 2Johns Hopkins University, 3Stevens Institute of Technology, 4Columbia University {chengq9, hengji}@illinois.edu
 1 Introduction
-Building robust language model (LM) agents for planning and reasoning has always been challenging.
-Recent efforts have explored how agents could compress and utilize memory (Wang et al., 2023a; Hu et al., 2023; Liu et al., 2023b; Liang et al., 2023b; Wang et al., 2024d, c; Zhong et al., 2024), perform complex reasoning (Wei et al., 2022; Kojima et al., 2022; Zhou et al., 2023a; Lin et al., 2024; Yao et al., 2023), planning (Wang et al., 2023b; Liu et al., 2023a; Hao et al., 2023; Yao et al., 2024; Zhou et al., 2024a), and reflection (Madaan et al., 2024; Zhang et al., 2024a, b; Miao et al., 2024; Dhuliawala et al., 2024) to improve task success rate.
-Integrating these capabilities, recent lines of work begin to build agents for embodied actions (Zheng et al., 2024; Huang et al., 2024a; Zhu et al., 2023) and tool use  (Schick et al., 2023; Qin et al., 2023; Qian et al., 2024) grounded in environments including the Web (Nakano et al., 2021; Furuta et al., 2024; Gur et al., 2024), games (Guo et al., 2023; Xu et al., 2023; Hu et al., 2024), and society (Park et al., 2023; Liu et al., 2023c; Li et al., 2023; Ren et al., 2024).
+Building robust language model (LM) agents to perform planning and reasoning has always been a challenging task.
+Recent efforts have explored how agents could compress and utilize memory (Wang et al., 2023a; Hu et al., 2023; Liu et al., 2023b; Liang et al., 2023b; Wang et al., 2024c; Zhong et al., 2024), perform complex reasoning (Wei et al., 2022; Kojima et al., 2022; Zhou et al., 2023a; Lin et al., 2024; Yao et al., 2023), planning (Wang et al., 2023b; Liu et al., 2023a; Hao et al., 2023; Yao et al., 2024; Zhou et al., 2024a), and reflection (Madaan et al., 2024; Zhang et al., 2024a, b; Miao et al., 2024; Dhuliawala et al., 2024) to improve task success rate.
+Integrating these capabilities, recent lines of work begin to build agents for embodied actions (Zheng et al., 2024; Huang et al., 2024a; Zhu et al., 2023) and tool use  (Schick et al., 2023; Qin et al., 2023; Qian et al., 2024; Wang et al., 2025; Qian et al., 2025a) grounded in environments including the Web (Nakano et al., 2021; Furuta et al., 2024; Gur et al., 2024), games (Guo et al., 2023; Xu et al., 2023; Hu et al., 2024), and society (Park et al., 2023; Liu et al., 2023c; Li et al., 2023; Ren et al., 2024).
 Figure 1: An agent with creative thinking should adapt its observation (e.g. hard texture of wood stick) into a novel tool-use strategy (e.g. prying objects open).
-The surge of agent systems also accelerates the development of simulation environments to evaluate an agent’s capabilities, including tasks like computer-based operations (Yao et al., 2022; Deng et al., 2024; Zhou et al., 2024b; Xie et al., 2024; Liu et al., 2024b), scientific research (Wang et al., 2022; Bran et al., 2023; Boiko et al., 2023; Huang et al., 2023a), and interactive experiences in text-based (Côté et al., 2019; Urbanek et al., 2019; O’Gara, 2023; Wu et al., 2024) or virtual sandbox game environments (Lin et al., 2023; BAAI, 2023; Wang et al., 2024a).
-However, most existing benchmarks are goal-oriented, emphasizing models’ planning, reasoning, and error-handling abilities, while overlooking an agent’s creativity: the capacity to think innovatively and adapt its observations to new, emergent, and unstructured scenarios.
+The surge of LM agent systems also accelerates the development of simulation environments, including tasks like computer-based operations (Yao et al., 2022; Deng et al., 2024; Zhou et al., 2024b; Xie et al., 2024; Liu et al., 2024b), scientific research (Wang et al., 2022; Bran et al., 2023; Boiko et al., 2023; Huang et al., 2023a), and interactive experiences in text-based (Côté et al., 2019; Urbanek et al., 2019; O’Gara, 2023; Wu et al., 2024) or virtual sandbox game environments (Lin et al., 2023; BAAI, 2023; Wang et al., 2024a).
+However, most existing benchmarks are usually goal-oriented, emphasizing models’ planning, reasoning, and error-handling abilities, while overlooking their creativity: the capacity to think innovatively and adapt their observations to new, unstructured scenarios.
 Current agents still significantly lack creativity in novel tool use (Zhang et al., 2023), as their training predominantly focuses on memorizing tool-task associations.
-This emphasis overshadows their ability to explore tool affordances and adapt to unstructured scenarios (Zhang et al., 2024c), both of which are critical for innovation.
-In robotics and autonomous systems, this limitation prevents agents from leveraging tools in unexpected ways to tackle unforeseen challenges (Bossema et al., 2023; Sandoval et al., 2022).
-In fact, creativity is widely recognized as a crucial component of intelligence.
+This emphasis overshadows their ability to explore tool affordances and adapt to unstructured scenarios (Zhang et al., 2024c).
+Despite this, creativity is still widely recognized as a crucial component of intelligence.
 In cognitive science, the well-established Triarchic Theory of Intelligence (Sternberg, 1984) divides intelligence into three components: practical, analytical, and creative.
-While current reasoning benchmarks primarily assess analytical intelligence through problem-solving, and simulation environments focus on practical intelligence by testing the application of knowledge in real-world scenarios, creative intelligence remains largely unaddressed.
-This gap is notable given its importance for enabling agents to demonstrate human-like adaptability and think beyond predefined parameters.
+While current reasoning benchmarks primarily assess analytical intelligence through problem-solving, and simulation environments focus on practical intelligence by testing knowledge application in real-world scenarios, creative intelligence remains largely unaddressed.
 To bridge this gap, we introduce EscapeBench, a benchmark that evaluates LM’s creative reasoning using scenarios inspired by room escape games.
-These scenarios test whether a model can escape a room through trial and error by solving unconventional puzzles that require “thinking outside the box” skills including creative observation, innovative tool use, and strategic problem-solving.
+These scenarios challenge conventional thinking through unusual settings and require “thinking outside the box” skills including creative tool use and strategic problem-solving.
 As shown in Figure 1, a wooden stick, typically used for walking or poking, has to be repurposed to pry open a lid due to its hard texture.
 This demands agents to perform adaptive reasoning under customized constraints.
-Overall, our benchmark sets itself apart through several distinctive features:
+Overall, our benchmark has several distinctive features:
   • Creative Tool Use: The tools at hand might be repurposed for creative use in order to solve the puzzle.
@@
 A human player may take up to an hour to complete one game.
-We benchmarked multiple open- and closed-source models within the BaseAgent framework, which incorporates working memory and Chain-of-Thought reasoning (Wei et al., 2022), by objectively evaluating their performance based on the total steps and hints used to complete each game.
+We benchmarked multiple models within the BaseAgent framework, which incorporates working memory and Chain-of-Thought reasoning (Wei et al., 2022).
 Our results show that even the best models struggle to complete the easiest game setting without hints, often requiring up to ten times the optimal steps and falling far behind human performance.
-These findings highlight how models tend to be constrained by conventional thinking patterns, struggling to break free from them and show creativity.
-To address this, we further introduce EscapeAgent, enhanced with Foresight for creative tool use and Reflection for implicit goal identification.
-The Foresight module encourages the agent to propose and self-evaluate creative hypotheses and predict outcomes for tool use before taking action, while the Reflection module helps the agent maintain an unsolved task list, providing dynamic reference to enhance its awareness of goals in future actions.
-Experiments show that EscapeAgent reduces hint reliance by nearly 50% and significantly lowers the number of action steps required to complete the game.
-It performs robustly across different difficulty levels, making progress more efficiently and exploring creative strategies with a higher success rate.
-In summary:
+These findings highlight how models tend to be constrained by conventional thinking patterns, struggling to break free and show creativity.
+To overcome this limitation, we introduce EscapeAgent, enhanced with Foresight for creative tool use and Reflection for implicit goal identification.
+Foresight enables the agent to propose and evaluate tool-use hypotheses before acting, while Reflection maintains an unsolved task list to guide future actions.
+Experiments show EscapeAgent reduces hint reliance by nearly 50%, lowers total action steps, and performs robustly across difficulty levels, achieving more efficient progress and higher action success rates with creative strategies.
+Our contributions include the following:
   • We identify challenges in LLM agent creative intelligence and introduce EscapeBench, a robust environment for evaluating agent creativity.
@@
 (2017).
-LMs have demonstrated notable creative capabilities across domains - they excel at generating narratives and poetry (Brown et al., 2020; Akoury et al., 2020), show effectiveness in tool creation and design (Qian et al., 2023; Cai et al., 2024), and augment human creativity through interactive ideation (Mialon et al., 2023).
+LMs have demonstrated notable creative capabilities across domains - they excel at generating narratives and poetry (Brown et al., 2020; Akoury et al., 2020), show effectiveness in tool creation and design (Qian et al., 2023; Cai et al., 2024), model real-world challenges (Qian et al., 2025b), and augment human creativity through interactive ideation (Mialon et al., 2023).
 In scientific discovery, research has also found that LM-generated ideas tend to be more novel but slightly less feasible than those from human experts (Si et al., 2024; Wang et al., 2024b).
@@
 Agent Evaluation in Simulated Environment.
-Existing work in agent evaluation primarily focuses on text-based and sandbox-like environments that assess an agent’s cognitive and behavioral abilities in goal-oriented tasks Zhou et al.
+Agent evaluation often focuses on text-based or sandbox environments for assessing cognitive and behavioral abilities in goal-oriented tasks Zhou et al.
 (2023b); Chen et al.
@@
 (2024); Deng et al.
-(2024), with emerging research exploring the integration of LM/VLM-enabled agents in robotics to address challenges in real-world, embodied settings Liang et al.
+(2024), with emerging work exploring LM/VLM-enabled agents in robotics for real-world challenges Liang et al.
 (2023a); Huang et al.
 (2023b, 2024b); Rana et al.
+(2023); Zhu et al.
+(2025); Yang et al.
 (2025).
@@
 (2023).
-Evaluation of agents in these environments typically relies on task-specific, cognitive, behavioral, and learning metrics to assess their ability to achieve objectives, reason adaptively, and collaborate.
-However, existing benchmarks often neglect creative adaptation in unfamiliar environments requiring unconventional tool use and proactive problem-solving.
-To address this gap, we introduce EscapeBench which evaluates agents’ creative reasoning and strategic problem-solving through navigating uncertain goal-achieving pathways, offering a new dimension of agent evaluation.
-Figure 2: An illustration of Scenes, Tools, and Items in the game engine design, and their relations with the model agent’s action space.
-Tools can be collected and used to craft new tools or applied to other items, while an item typically requires a click, input, or tool application to trigger specific effects.
+The settings typically rely on task-specific metrics for goal achievement but overlook creative and proactive problem-solving in unfamiliar contexts.
+To address this, we introduce EscapeBench to evaluate agents’ creative reasoning in navigating uncertain goal pathways, offering a novel approach to agent assessment.
+Figure 2: An illustration of Scenes, Tools, and Items in the game and their relations with agent action space.
+Tools can be collected for “Apply” and “Craft”, while items require “Input”, “Click” or “Apply” of tools to trigger effects.
 3 EscapeBench Construction
-3.1 Motivation
-Most current agent benchmarks emphasize explicit, goal-oriented tasks, where decisions rely on commonsense knowledge encoded within the agent model’s parametric space.
-For instance, a typical task in AgentBench (Liu et al., 2024b) might ask: “What musical instruments did the Minnesota-born Nobel prize winner play?” Using the Freebase search API, the agent can chart a clear pathway to the goal, identifying relevant data and filtering out irrelevant results.
-This process is deterministic and fact-based, with each step reflecting the agent’s analytical skill in processing search outputs.
-While such benchmarks effectively test an agent’s information-seeking and retrieval capabilities, they often overlook its ability to engage in creative problem-solving—an area that extends beyond analytical and practical intelligence.
-This leads to our core research question: How to build an environment that benchmarks an agent’s creative intelligence?
-To answer this, an agent must demonstrate out-of-the-box thinking in its reasoning.
-Since the primary function of agents lies in their tool-use abilities, we aim to design a benchmarking environment that emphasizes creative tool use.
-Room escape game scenarios, by nature, are well-suited for this purpose, as they require a player to creatively make use of all the things at hand to solve difficult puzzles.
-In the following sections, we introduce the components of our game engine (Section 3.2), the action space available to agents (Section 3.3), the data annotation process (Section 3.4), and preliminary case study results (Section 3.5).
+Most agent benchmarks focus on explicit, goal-oriented tasks grounded in commonsense knowledge, where agents can chart clear pathways to achieve goals using analytical and practical intelligence, but they often overlook creative intelligence.
+This raises our core research question: How to build an environment that benchmarks an agent’s creative intelligence?
+Given that tool use is central to agent functionality, we propose room escape game scenarios, which naturally require creative tool use to solve complex puzzles, as an ideal environment for this evaluation.
 3.1 Engine Design
-Our engine aims to simulate the room escape environment that i) receives agent actions and ii) makes corresponding environment responses as agent action’s reward.
-It builds up a logical skeleton that ensures a smooth benchmarking process once the game configuration is given.
+Our game engine aims to simulate the room escape environment that i) receives agent actions and ii) makes corresponding environment feedback as agent action’s reward.
 Specifically, our game engine involves three key components:
@@
   • Tools: Objects that could be collected in each scene, usually applied to other items to take effect or to other tools to craft new ones.
-The interaction of these three key components forms a game scenario’s basic logic.
-For the game engine example in Figure 2, different scenes inside a van connect with each other forming a graph to represent their physical connectivity via connectors, such as doors or tunnels.
-The tool key chain collected in a scene will be placed in the bag for future use, while the item wire iron is waiting for something sharp to cut it off to trigger some effects.
+The interaction of these components defines a game’s basic logic.
+In the van example from Figure 2, scenes are connected into a graph, representing physical connectivity via doors or tunnels.
+The tool key chain is collected in the bag for future use, while the wire iron awaits something sharp to cut it to trigger effects.
 Please refer to Appendix A for more detailed examples and explanations.
 3.2 Action Space
-The model agent could take five different actions in each scene.
-While the action space is well-defined, the parameter space—regarding the scenes, items, or tools involved in these actions—is high-dimensional, thus enabling dynamic interactions.
+The model agent could take five different actions.
+While the action space is well-defined, the parameter space—regarding the scenes, items, or tools involved in these actions—is high-dimensional, thus allowing for dynamic interactions.
   • Move (Scene): Move to an adjacent scene.
@@
   • Craft (Tool, Tool): Use two tools in the bag to craft a new one.
-Figure 2 illustrates the connections between the components of the game engine and the model agent’s action space.
+Figure 2 illustrates the connections between game engine components and agent action space.
 Among all the actions, “Apply” and “Craft” stand out as the most creativity-driven, as they require the agent to think innovatively about how to use or craft tools in an unseen way during its training.
 We delve into specific examples in Section 3.4.
+Figure 3: Statistics of total Scenes, Tools, and Items across all game settings. “Key Steps” refer to the essential bottleneck actions required to complete the game.
 Table 1: Rules and examples for different difficulty levels.
@@
 represents the game engine’s feedback when an unexpected action targets an Item or Tool.
-Figure 3: Statistics of total Scenes, Tools, and Items across all game settings. “Key Steps” refer to the essential bottleneck actions required to complete the game.
 3.3 Annotation and Statistics
-Building on existing online room escape games and puzzle-solving logic [^2] https://spotlight.ee, we present EscapeBench, featuring 12 scenarios and 36 settings across three difficulty levels.
-Each scenario, such as a laundry room or basement exploration, is manually annotated for high-quality game engine performance.
-Difficulty levels are defined by the clarity of descriptions and the feedback provided for unexpected actions, detailed in Table 1.
+Building on existing online room escape games and puzzle-solving logic [^2] https://spotlight.ee, we introduce EscapeBench, featuring 36 game settings spanning three difficulty levels.
+Each scenario includes three versions with consistent solution logic but varying in description clarity and feedback granularity.
+A detailed example is shown in Table 1.
+All scenes, items, and tools are manually annotated to ensure high quality.
 These scenarios emphasize creative tool use and crafting strategies, challenging agents throughout the game, making EscapeBench a robust environment for testing creativity.
-We make all our scenarios text-based to accommodate the evaluation of top-performing language models.
-We present in Figure 3 detailed statistics on total scenes, tools, items, and key steps involved in all game settings.
+A detailed statistic is presented in Figure 3.
 Table 2: Creative reasoning cases in EscapeBench and model’s responses versus actual creative uses of tools.
 3.4 Preliminary Study
-To verify the creative reasoning required in our designed games and assess the current LMs’ ability to perform such reasoning, we conduct a preliminary case study in Table 2.
-We sample diverse real scenarios from EscapeBench’s game design and test them on closed-source GPT-4o (Hurst et al., 2024) and open-source LLama-3.1-70B (Dubey et al., 2024).
-For each model, we instruct it to think creatively about what it might achieve with the provided Items and Tools in room escape scenarios.
-Our results reveal that: i) EscapeBench offers a wide range of creative reasoning challenges, encompassing unconventional tool usage, implicit numerical puzzle solving, innovative tool crafting strategies, etc.
-ii) Both closed- and open-source models demonstrate significant limitations in creativity and imagination, particularly in identifying implicit goals and devising creative strategies.
-These findings emphasize the complexity and necessity of our benchmark as well as the current gaps in models’ creativity.
+We sample scenarios from EscapeBench and test GPT-4o (Hurst et al., 2024) and LLama-3.1-70B (Dubey et al., 2024)’s creative reasoning performance through case studies in Table 2.
+Our results reveal that: i) EscapeBench presents diverse creative reasoning challenges, including unconventional tool use, implicit numerical puzzles, and innovative tool crafting.
+ii) Both closed- and open-source models struggle with creativity, especially in identifying implicit goals and forming creative strategies.
+These findings highlight the complexity of EscapeBench and the gaps in model creativity.
 Figure 4: Illustration of the EscapeAgent design.
@@
 4 EscapeAgent Design
-Our preliminary study highlights that current LMs often struggle with identifying implicit goals and engaging in the creative reasoning needed to solve complex puzzles.
-To overcome these challenges and maximize the LM’s potential, we present EscapeAgent, a novel framework designed to uncover tasks dynamically and reason creatively about tool use.
-EscapeAgent addresses two core challenges identified by EscapeBench:
-  • Uncertain Goal Pathways: We introduce a Reflection module, which dynamically maintains a task list by adding, updating, or removing tasks after each trial-and-error action.
-This approach fosters proactive task discovery and sharpens the agent’s focus on specific goals (Section 4.2).
-  • Creative Tool Use: We design a Foresight module that facilitates explicit reasoning about creative tool applications and their alignment with specific tasks.
-It enables the agent to hypothesize and evaluate strategies before execution, thus promoting problem-solving (Section 4.3).
-By integrating both modules with a BaseAgent, EscapeAgent excels at managing Super-long Reasoning Chains, effectively performing over 1,000 consecutive actions while maintaining logical coherence.
-This robust design greatly enhances the language model’s capabilities in creativity, problem-solving, and strategic thinking.
+To address challenges identified in the preliminary study, we introduce EscapeAgent, a framework addressing two core issues from EscapeBench:
+  • Uncertain Goal Pathways: A Reflection module dynamically manages a task list, refining goals through trial-and-error to enhance action focus and proactive task discovery (Section 4.2).
+  • Creative Tool Use: A Foresight module enabling explicit reasoning about tool applications, allowing the agent to hypothesize and evaluate strategies before execution (Section 4.3).
+Integrating both modules with a BaseAgent, EscapeAgent excels in handling Super-long Reasoning Chains and significantly boosts the model’s creativity, problem-solving, and strategic thinking.
 4.1 BaseAgent
@@
 4.2 Reflection Module
-The Reflection Module enables the agent to maintain a structured task list, updated through:
+The Reflection Module manages a structured task list updated through three actions:
   • New: Add a newly identified, unsolved task.
@@
   • Delete: Remove a task when its goal is achieved.
-Each task list entry includes the task name, the target item, and previously failed actions. This organization prevents the agent from repeating ineffective actions, thus promoting efficiency.
-The reflection module is triggered after every non-move action, using environment feedback to guide task list updates.
-For example, in Figure 4 (right), Task 1 is deleted once the machine successfully starts.
-By encouraging targeted problem-solving rather than random exploration, the reflection module significantly improves task-solving efficiency.
-See Section C.2 for more details.
+Each entry includes the task name, target item, and failed actions, preventing repeated mistakes and improving efficiency.
+Triggered after non-move actions, it uses feedback to update the task list.
+For example, in Figure 4 (right), Task 1 is deleted once the machine starts, encouraging focused problem-solving over random exploration.
+See Section C.2 for details.
 4.3 Foresight Module
 The Foresight Module enhances creative reasoning by explicitly evaluating tool use and problem-solving strategies. It activates in two cases:
-  • When a new task is identified: The agent evaluates existing tools it has and hypothesizes potential actions to achieve the task.
-  • When a new tool is collected: The agent assesses how the tool might contribute to solving existing tasks or crafting new tools.
-If the module generates at least one valid action hypothesis, the agent enters a “Try Action” state, where it systematically tests the proposed actions in subsequent steps.
-Otherwise, it remains in a “Free Explore” state, operating like the BaseAgent to explore the environment.
-For instance, in Figure 4, the foresight module identifies the action of clicking a button on the machine, thus guiding the agent into the “Try Action” state.
-Overall, this module enables the agent to adapt flexibly under customized scenarios, make bold hypotheses, and execute targeted action trials efficiently.
-See Section C.3 for further details.
-5 Experiments
-We divide our experiments into two parts: i) Benchmarking the creativity of current models using the BaseAgent (Section 5.2), and ii) Demonstrating the effectiveness of EscapeAgent (Section 5.3).
-5.1 Settings
-Environment.
-Experiments are conducted on 36 game settings.
-An agent is considered to be making progress if it either achieves a key step (bottleneck actions defined in Figure 3) or collects a new tool.
-Agents will receive help if they fail to make progress for 50 consecutive actions (see Section D.1), thus ensuring full completion of the game.
-The working memory length is set to 10 due to context window limitations.
-Models.
-We evaluate both closed- and open-source models: GPT-4o, GPT-4o-mini (Hurst et al., 2024), Claude-3.5 (Anthropic, 2024), Gemini-1.5 (Team et al., 2024), Llama-3.1 (Dubey et al., 2024), Qwen-2.5 (Team, 2024), DeepSeek-LLM (Liu et al., 2024a), Phi-3.5 (Abdin et al., 2024), Yi (Young et al., 2024), Ministral (MistralAI, 2024).
-Models with fewer than 7B parameters are excluded due to near-random behavior.
-For consistency, we set the sampling temperature to $T=0$ and $n=1$.
-Metrics. We use two main metrics including:
-  • Hints Used: Total hints used in a game.
-  • Total Steps: Total actions taken in a game.
-Auxiliary metrics for analysis include:
-  • Early Exit Progress: Proportion of key steps and tools collected before needing a hint (game progress before needing a hint for the first time).
-  • Tool Hints Used (percentage): Hints used for tool collection (normalized by total tools).
-  • Key Step Hints Used (percentage): Hints used for key steps (normalized by total key steps).
-Results are micro-averaged across the 36 settings.
+  • New Task Identified: The agent hypothesizes potential actions to achieve it using available tools.
+  • New Tool Collected: The agent assesses its use for solving existing tasks or crafting new tools.
+If a valid hypothesis is proposed, the agent enters “Try Action” state to test it; otherwise, it stays in “Free Explore” state, operating like the BaseAgent.
+For example in Figure 4 (left), the agent identifies clicking the button as an action worth trying, thus guiding it into the “Try Action” state.
+It enables the agent to adapt flexibly under customized scenarios, make bold hypotheses, and execute targeted trials efficiently.
+See Section C.2 for details.
 Model Name Hints Used Total Steps Early Exit  Progress (%) Tool Hints Used (percentage) Key Steps Hints Used (percentage)
@@
 Average Human 4.33 257.83 59.65 0.17 (0.69%) 4.17 (12.28%)
-Oracle Chain (Upper Bound) 0.00 107.83 100.00 0.00 0.00
-Table 3: Benchmarking results of baseline BaseAgent with different core models on EscapeBench.
+Table 3: Benchmarking results of BaseAgent with different core models on EscapeBench.
+An oracle action chain’s total step is only 107.83 on average.
 Both closed- and open-source models rely heavily on hints to complete the escape compared to human performance, with smaller-scale models exhibiting a particularly high dependency.
-5.2 Benchmarking Results
-We benchmark current models using the BaseAgent framework, with results in Table 3 showing that large-scale closed-source models, such as GPT-4o, Claude-3.5, and Gemini-1.5, consistently outperform smaller models.
-Key insights include:
-  • Most hints are used on key steps, which require creative reasoning, while models can often collect tools through random exploration.
-  • Models require significantly more action steps and hints than the average human player, with some using up to  20x more steps than the most efficient action chain.
-Figure 5: Ablation of difficulty through Hints Used. Harder game settings generally require more hints to complete, and vice versa.
-Figure 6: Ablation of difficulty through Total Steps. Harder game settings generally require more total steps to complete, and vice versa.
-Complex environment causes challenges.
-We show in Figure 5 and Figure 6 that, generally, harder game settings require more steps and hints for an agent to solve.
-Since our difficulty setting depends solely on the granularity and usefulness of descriptions and feedback (see Table 1), our results demonstrate that the way the environment is presented can impact difficulty, even when the core game logic remains unchanged.
+5 Experiments
+We divide experiments into: i) Benchmarking model creativity within the BaseAgent, and ii) Evaluating EscapeAgent’s effectiveness.
+5.1 Settings
+Environment.
+Experiments are conducted on 36 game settings.
+An agent is considered to be making progress if it either achieves a key step (defined in Figure 3) or collects a new tool.
+Agents will receive help if they fail to make progress for 50 consecutive actions (see Section D.1), thus ensuring full completion of the game.
+The working memory length is set to 10.
+Models.
+We evaluate both closed- and open-source models: GPT-4o, GPT-4o-mini (Hurst et al., 2024), Claude-3.5 (Anthropic, 2024), Gemini-1.5 (Team et al., 2024), Llama-3.1 (Dubey et al., 2024), Qwen-2.5 (Team, 2024), DeepSeek-LLM (Liu et al., 2024a), Phi-3.5 (Abdin et al., 2024), Yi (Young et al., 2024), Ministral (MistralAI, 2024).
+Models with fewer than 7B parameters are excluded due to near-random behavior.
+For consistency, we set sampling temperature to $T=0$ and $n=1$.
+Metrics.
+We use two main metrics including:
+  • Hints Used: Total hints used in a game.
+  • Total Steps: Total actions taken in a game.
+Auxiliary metrics for analysis include:
+  • Early Exit Progress: Proportion of key steps and tools collected before needing a hint (game progress before needing a hint for the first time).
+  • Tool Hints Used (percentage): Hints used for tool collection (normalized by total tools).
+  • Key Step Hints Used (percentage): Hints used for key steps (normalized by total key steps).
+Results are micro-averaged across the 36 settings.
 Figure 5: Distribution of Key Steps Hints Used, categorized by different actions.
 Colored bars represent the percentage of hints used for each action type relative to the total key steps for that type (See right of Figure 3).
-“Craft” and “Input” account for the highest percentage of relative hint usage, highlighting their challenge.
-Input and craft are the most challenging.
-As shown in Figure 7, while “Apply” actions require the most hints overall, “Input” and “Craft” actions show the highest relative hint usage compared to their total occurrence in key steps.
-This reflects the “Input” action’s large parametric space, making random guesses impractical.
-Additionally, the implicit numerical challenges presented by “Input” often require synthesizing multiple pieces of evidence, adding to their difficulty.
-The results also suggest that the model demonstrates a degree of creativity, enabling effective “Apply” of tools, but “Craft” actions are inherently more creativity-demanding.
-Table 4: Error Analysis of BaseAgent’s inefficiency or failures.
-All cases are selected in the same game setting.
-Superficial attempts dominate larger-scale models, inspiring EscapeAgent’s design to address them.
+Table 4: Error Analysis of BaseAgent’s inefficiency or failures. All cases are selected in the same game setting.
+Model Name Hints Used Total Steps Early Exit  Progress (%) Tool Hints Used (percentage) Key Steps Hints Used (percentage)
+GPT-4o $5.03_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 5.27}}$ $452.75_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 270.86}}$ $47.03_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 22.28}}$ $0.33_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 1.84}}$ (1.19%) $4.70_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 3.44}}$ (13.74%)
+GPT-4o-mini $10.58_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 4.61}}$ $752.25_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 250.14}}$ $28.17_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 12.11}}$ $1.14_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 0.86}}$ (4.16%) $9.44_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 3.75}}$ (28.53%)
+Llama-3.1-70B-Instruct $7.92_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 6.61}}$ $645.19_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 337.23}}$ $31.44_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 12.44}}$ $1.42_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 1.69}}$ (5.44%) $6.56_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 4.86}}$ (19.15%)
+Qwen2.5-72B-Instruct $9.72_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 6.78}}$ $746.61_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 355.89}}$ $28.62_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 16.16}}$ $2.72_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 2.61}}$ (10.61%) $7.00_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 4.17}}$ (20.71%)
+DeepSeek-LLM-67b-Chat $20.14_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 5.36}}$ $1285.03_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 273.44}}$ $15.30_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 8.67}}$ $10.81_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 0.31}}$ (42.72%) $9.34_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 5.66}}$ (27.20%)
+Yi-1.5-34B-Chat $22.59_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 1.41}}$ $1468.03_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 105.30}}$ $12.04_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 0.08}}$ $10.42_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 2.31}}$ (41.61%) $12.19_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 3.73}}$ (35.71%)
+Phi-3-medium-128k-instruct $25.75_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 6.44}}$ $1513.69_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 357.50}}$ $9.99_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 2.65}}$ $11.20_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 0.91}}$ (44.99%) $14.55_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 5.56}}$ (43.35%)
+Llama-3.1-8B-Instruct $19.81_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 6.05}}$ $1271.53_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 271.77}}$ $19.22_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 9.12}}$ $6.64_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 0.17}}$ (25.23%) $13.22_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 5.89}}$ (39.34%)
+Ministral-8B-Instruct $19.47_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 5.84}}$ $1233.72_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 323.25}}$ $19.34_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 10.37}}$ $6.61_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 0.56}}$ (25.56%) $12.86_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 5.33}}$ (38.93%)
+Qwen2.5-7B-Instruct $27.53_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 4.67}}$ $1639.58_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 310.84}}$ $8.56_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 2.04}}$ $13.00_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 0.81}}$ (52.66%) $14.58_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 3.89}}$ (43.99%)
+Table 5: Benchmarking Results of EscapeAgent with different core models.
+Nearly all the performance raises compared to Table 3, showcasing the effectiveness of EscapAgent in promoting the agent’s creativity.
+5.2 Benchmarking Results
+We benchmark current models using the BaseAgent framework, with results in Table 3 showing that large-scale closed-source models consistently outperform smaller models. Key insights include:
+  • Most hints are used on key steps, which demand creative reasoning, while models may often collect tools through random exploration.
+  • Models require significantly more action steps and hints than the average human, and up to 2̃0x more steps than the most efficient action chain.
+We further present an ablation study of total hints and steps used in Section D.3.
+Input and craft are the most challenging actions.
+As shown in Figure 5, while “Apply” actions require the most hints in absolute terms, “Input” and “Craft” actions have the highest relative hint usage compared to the total number of key steps.
+This likely reflects the large parametric space of “Input” actions, where random guesses are impractical, and the creativity-demanding nature of “Craft” actions.
 Error analysis.
@@
 Our EscapeAgent addresses this by promoting more purposeful and creative actions.
-Model Name Hints Used Total Steps Early Exit  Progress (%) Tool Hints Used (percentage) Key Steps Hints Used (percentage)
-GPT-4o $5.03_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 5.27}}$ $452.75_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 270.86}}$ $47.03_{{\color[rgb]{1,0,0}\uparrow 22.28}}$ $0.33_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 1.84}}$ (1.19%) $4.70_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 3.44}}$ (13.74%)
-GPT-4o-mini $10.58_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 4.61}}$ $752.25_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 250.14}}$ $28.17_{{\color[rgb]{1,0,0}\uparrow 12.11}}$ $1.14_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 0.86}}$ (4.16%) $9.44_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 3.75}}$ (28.53%)
-Llama-3.1-70B-Instruct $7.92_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 6.61}}$ $645.19_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 337.23}}$ $31.44_{{\color[rgb]{1,0,0}\uparrow 12.44}}$ $1.42_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 1.69}}$ (5.44%) $6.56_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 4.86}}$ (19.15%)
-Qwen2.5-72B-Instruct $9.72_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 6.78}}$ $746.61_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 355.89}}$ $28.62_{{\color[rgb]{1,0,0}\uparrow 16.16}}$ $2.72_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 2.61}}$ (10.61%) $7.00_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 4.17}}$ (20.71%)
-DeepSeek-LLM-67b-Chat $20.14_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 5.36}}$ $1285.03_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 273.44}}$ $15.30_{{\color[rgb]{1,0,0}\uparrow 8.67}}$ $10.81_{{\color[rgb]{1,0,0}\uparrow 0.31}}$ (42.72%) $9.34_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 5.66}}$ (27.20%)
-Yi-1.5-34B-Chat $22.59_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 1.41}}$ $1468.03_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 105.30}}$ $12.04_{{\color[rgb]{1,0,0}\uparrow 0.08}}$ $10.42_{{\color[rgb]{1,0,0}\uparrow 2.31}}$ (41.61%) $12.19_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 3.73}}$ (35.71%)
-Phi-3-medium-128k-instruct $25.75_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 6.44}}$ $1513.69_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 357.50}}$ $9.99_{{\color[rgb]{1,0,0}\uparrow 2.65}}$ $11.20_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 0.91}}$ (44.99%) $14.55_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 5.56}}$ (43.35%)
-Llama-3.1-8B-Instruct $19.81_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 6.05}}$ $1271.53_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 271.77}}$ $19.22_{{\color[rgb]{1,0,0}\uparrow 9.12}}$ $6.64_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 0.17}}$ (25.23%) $13.22_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 5.89}}$ (39.34%)
-Ministral-8B-Instruct $19.47_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 5.84}}$ $1233.72_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 323.25}}$ $19.34_{{\color[rgb]{1,0,0}\uparrow 10.37}}$ $6.61_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 0.56}}$ (25.56%) $12.86_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 5.33}}$ (38.93%)
-Qwen2.5-7B-Instruct $27.53_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 4.67}}$ $1639.58_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 310.84}}$ $8.56_{{\color[rgb]{1,0,0}\uparrow 2.04}}$ $13.00_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 0.81}}$ (52.66%) $14.58_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 3.89}}$ (43.99%)
-Table 5: Benchmarking Results of our approach EscapeAgent with different core models.
-Nearly all the performance raises compared to Table 3, showcasing the effectiveness of EscapAgent in promoting the agent’s creativity.
 5.3 EscapeAgent Results
-The introduction of the Reflection and Foresight modules in EscapeAgent, as shown in Table 5, leads to a marked reduction in hint usage and total steps, with large-scale models benefiting the most.
-Key insights include:
-  • Larger models like GPT-4o consistently outperform smaller models, suggesting that while the new modules aid creative reasoning, the core model’s capabilities remain a crucial factor.
-  • Early exit progress improves significantly across models, despite the exponentially increasing difficulty of maintaining consecutive meaningful progress without hints.
-Notably, GPT-4o achieves near-human-level performance, demonstrating the effectiveness of the enhanced framework.
-Figure 8: Distribution of step intervals for progress made through Tool Collection and Key Step achievement. EscapeAgent uses fewer steps to achieve the next progress and relies less on hints.
+The introduction of the Reflection and Foresight in EscapeAgent, as shown in Table 5, significantly reduces hint uses and total steps, with larger models benefiting the most. Key insights include:
+  • Larger models still outperform smaller ones, suggesting while new modules aid creative reasoning, the core model’s capabilities remain crucial.
+  • Early exit progress improves across models, despite the exponentially increasing difficulty of consecutively making progress without hints, demonstrating EscapeAgent’s effectiveness.
 Figure 6: The accumulated number of completed games (in total 36) relative to total steps a game setting takes.
 EscapeAgent, shown in dotted lines, completes games in fewer steps, demonstrating greater efficiency.
-EscapeAgent progresses more efficiently.
-Figure 9 demonstrates EscapeAgent’s more efficient progress, as the slopes of the dotted lines are consistently steeper.
-We further show in Figure 8 that EscapeAgent requires fewer actions to achieve the next valid Key Steps compared to BaseAgent, indicating better creative reasoning.
-The spike at 50 steps corresponds to explicit hints provided after the model’s prolonged inactivity for consecutively 50 steps, and the lower pink bar here further highlights EscapeAgent’s reduced hint dependency.
-Interestingly, for all models, sudden progress after 15 steps without hints is rare, highlighting a lack of spontaneous insights that humans often demonstrate, even in the EscapeAgent framework.
-Figure 10: A comparison of action trial times distribution for a specific item before success. EscapeAgent tries more actions with a higher overall success rate.
-EscapeAgent tries more with higher overall success.
-In Figure 10, we analyze the valid actions agents attempt on different items in scenes before successfully operating them.
-While the BaseAgent exhibits a higher first-attempt success rate, EscapeAgent achieves greater effectiveness by making multiple attempts, leading to a significantly higher overall success rate within 10 trials.
-This difference can be attributed to EscapeAgent’s strategy of proposing multiple viable actions simultaneously.
-Although this increases the likelihood of eventually succeeding by trying more actions, it does not prioritize the most-likely-to-succeed action first.
-As a result, the BaseAgent appears more efficient on its initial trial, but its superficial approach, as highlighted in Table 4, limits its overall performance.
-The EscapeAgent’s design effectively addresses this limitation by leveraging a more exploratory approach, which proves advantageous in complex scenarios requiring creativity.
 Figure 7: Case study on Human, BaseAgent, and EscapeAgent’s progress map corresponding to six game settings.
 Although EscapeAgent uses 40% fewer hints and makes significant progress independently, it still falls far short of average human performance, often requiring twice as many total steps to complete a game.
+Figure 8: Distribution of step intervals for progress made through tool collection and key step achievement. EscapeAgent uses fewer steps to achieve the next progress and relies less on hints.
+EscapeAgent progresses more efficiently.
+As shown in Figure 6, EscapeAgent demonstrates steeper progress slopes, reflecting greater efficiency.
+Figure 8 further shows that EscapeAgent requires fewer actions to reach the next key step, indicating stronger creative reasoning ability.
+The spike at 50 steps corresponds to hints provided after prolonged inactivity, so the lower pink bar here further highlights EscapeAgent’s reduced hint dependency.
+Notably, across all models, progress after 15 steps without hints is rare, underscoring a lack of spontaneous insights typically seen in humans.
 Case study.
-Figure 11 presents a case study of agent progress relative to action steps across six game settings.
-Red dots indicate steps where hints were provided.
-We observe that i) EscapeAgent requires significantly fewer hints and achieves a steeper progress curve; ii) EscapeAgent demonstrates the ability to make consecutive progress in shorter intervals (e.g., the light blue and pink lines for both GPT-4o and Llama-3.1-70B); iii) harder scenarios remain challenging, especially for BaseAgent, which often relies heavily on hints to make progress (e.g., the later part of the yellow line for GPT-4o); and iv) Average human still performs far better than all agents.
-Humans rarely make mistakes shown in Table 4, while both agents still struggle with short memory due to context length and creative tool use strategies.
-These findings emphasize the need for further improvements and highlight EscapeBench’s challenge to even the most advanced models.
+Figure 7 illustrates agent progress relative to action steps across six game settings, with red dots marking hint provided.
+We observe that: i) EscapeAgent requires fewer hints and achieves steeper progress; ii) it can make consecutive progress in shorter intervals; iii) harder scenarios remain challenging, especially for BaseAgent, which heavily relies on hints; iv) Average human performs far better.
+Humans rarely make mistakes shown in Table 4, while agents still struggle with short memory due to context length and creative tool use strategies.
+These emphasize the need for further improvements and highlight EscapeBench’s challenge to even the most advanced models.
+6 Further Analysis
+6.1 Evaluation across Diverse Models
+To enhance the diversity of model scales and domains in our evaluation, we present additional experimental results on smaller-scale and domain-specific models.
+These results supplement our main findings and provide further insights into the scalability and generalizability of our framework.
+In addition to the models presented in out main table, we also evaluated sub-7B models, including the latest Llama-3.2-1B and 3B variants.
+However, these models demonstrated extremely limited performance.
+In even the simplest game scenarios under normal difficulty, these models frequently resorted to random guessing, with over 95% of key actions requiring hint usage.
+The inefficiency translated into significantly inflated action sequences and minimal task completion, leading us to our conclusion that 7B is the minimum viable scale for meaningful evaluation in our benchmark.
+Model Hints Used Total Steps Early Exit Progress (%)
+Qwen-2.5-7B-Instruct 22.6 1329.8 5.93
+Qwen-2.5-Coder-7B 22.4 1339.8 12.99
+Qwen-2.5-Math-7B 45.4 2384.4 0.00
+Table 6: Performance of Domain-Specific Models on EscapeAgent Benchmark
+To examine domain specialization effects, we benchmarked two variants from the Qwen-2.5 model series, focused respectively on coding and mathematical tasks.
+As shown in Table 6, the Qwen-2.5-Coder-7B achieved performance close to the general-purpose 7B-Instruct model, indicating effective domain adaptation without significant loss of reasoning flexibility.
+In contrast, the Qwen-2.5-Math-7B model struggled, failing to initiate or complete key actions autonomously.
+This suggests that excessive specialization, such as strict math alignment, may hinder the model’s ability to generalize and engage in creative, multi-step tasks.
+6.2 Ablation Study on Key Modules
+To evaluate the individual contributions of the Foresight and Reflection modules in EscapeAgent, we conducted an ablation study using two modified agent variants:
+  • EscapeAgent (only Foresight): Disables the task list maintained by the Reflection module, allowing action proposals based solely on environmental observation.
+  • EscapeAgent (only Reflection): Disables the Foresight module, relying instead on the task list to guide decision-making.
+Model Hints Used Total Steps Early Exit Progress (%)
+GPT-4o, EscapeAgent 5.00 452.00 44.55
+GPT-4o, only Reflection 6.75 570.08 37.00
+GPT-4o, only Foresight 7.17 593.92 48.89
+GPT-4o, BaseAgent 9.83 707.33 19.85
+Llama-3.1-70B, EscapeAgent 7.17 624.67 28.31
+Llama-3.1-70B, only Reflection 11.33 878.67 28.78
+Llama-3.1-70B, only Foresight 11.00 877.25 22.29
+Llama-3.1-70B, BaseAgent 14.67 981.42 17.28
+Table 7: Ablation results on GPT-4o and Llama-3.1-70B-Instruct backbones.
+We benchmarked these variants alongside the full EscapeAgent and the BaseAgent across all game scenarios with normal difficulty.
+Across both GPT-4o and Llama-3.1-70B in Table 7, each module independently improves performance relative to the BaseAgent.
+The full EscapeAgent consistently achieves the best results, confirming that the Foresight and Reflection modules are complementary.
+Notably, Reflection alone often outperforms Foresight alone, likely because the Foresight module relies on the task list to avoid redundant or failed actions.
+These results support the conclusion that combining both modules is essential to fully realize EscapeAgent’s potential.
 7 Discussions and Future Directions
 LM’s creativity for benchmarking.
-Our experiments with EscapeBench reveal key insights into the capabilities and limitations of current language models in creative reasoning and tool use.
-Even the most advanced model, EscapAgent, requires more hints and takes twice as many steps as the average human to solve tasks.
-The benchmark demonstrates that these advanced models struggle with implicit goal identification and creative problem-solving, often relying heavily on extensive hints.
-This underscores a significant gap in the evaluation of AI agents: while analytical and practical intelligence are well-assessed, creative intelligence remains largely underexplored.
-Addressing this gap may require fundamentally improving how LMs organize knowledge, linking objects not just by words but by their affordances—how their properties and functions relate to potential uses.
+Our experiments reveal that even the most advanced language models within EscapeAgent require more hints and twice as many steps as the average human, exposing limitations in creative reasoning and tool use.
+The benchmark highlights that while analytical and practical intelligence is well-assessed, creative intelligence remains a critical gap.
+Addressing this gap may require enhancing LMs to link knowledge and objects through affordances—their properties and functions—to foster creativity.
+Theoretical Foundations for AI Creativity.
+Human creativity, characterized by generating novel ideas and adapting to complexity, arises from the interplay of stochastic neuronal noise and structured, learned information Dainys (2024); Malach (2024).
+In contrast, AI creativity relies on trained data patterns and algorithms.
+Boden identifies three mechanisms driving AI creativity: combining familiar ideas, exploring conceptual spaces, and enabling transformative innovations Boden (1998).
+Integrating insights from psychology and neuroscience may further enhance AI’s creativity.
 Human-AI Collaboration.
-Human-AI collaboration in escape room puzzles holds the potential to significantly enhance problem-solving efficiency and introduce a new paradigm of learning.
-Humans contribute intuitive insights and creative ideas that language models are unlikely to generate independently, while AI excels in systematic tasks such as data analysis and information aggregation.
-Furthermore, this collaboration leverages the strength of chain-of-thought reasoning, a process through which AI organizes and evaluates information in a logical, step-by-step manner.
-Integrating human creativity with AI’s structured problem-solving capabilities together in the future might not only foster innovative strategies but also enable deeper learning experiences, making the problem-solving process more dynamic and efficient.
-Theoretical Foundations for AI Creativity.
-Understanding the cognitive mechanisms behind human creativity is essential for designing AI systems that emulate or surpass human creative processes.
-Human creativity is a multifaceted phenomenon involving the generation of novel and valuable ideas, problem-solving, and adaptation to complex situations Dainys (2024).
-Neurologically, it arises from the interplay between stochastic neuronal noise and structured, learned information, driven by spontaneous brain activity fluctuations Malach (2024).
-This contrasts with AI’s reliance on data patterns and algorithmic processes.
-Boden identifies three AI creativity mechanisms: combining familiar ideas in novel ways, exploring conceptual spaces, and enabling transformative innovations Boden (1998).
-Our research shows that even advanced models like GPT-4o struggle with implicit goal identification and creative problem-solving, often requiring extensive prompting.
-The EscapeAgent framework significantly reduces this dependency and improves task-solving efficiency, indicating that these modules effectively overcome barriers to creative reasoning.
-However, the results also highlight the importance of the core model’s capabilities, as larger models like GPT-4o benefit more from the framework than smaller models.
-This suggests further research is needed to address current models’ limitations in generating truly novel ideas.
-Interdisciplinary approaches integrating psychology, neuroscience, and model architecture can advance agent creativity further.
-Multimodal Integration.
-Expanding the escape room environment to include multimodal data, such as visual and voice cues, offers a promising avenue for enhancing both the agent’s performance and the realism of the scenarios.
-While integrating vision-language models could enable agents to interpret visual clues more naturally, such an extension would also require robust visual understanding and reasoning capabilities to handle the complexity of tasks effectively.
-Additionally, incorporating multimodal interactions presents opportunities to study how agents synthesize information across modalities, such as correlating visual patterns with instructions or adapting strategies based on dynamic, multimodal feedback.
-Future work should explore how to seamlessly integrate and harmonize these diverse data types into the benchmark, pushing the boundaries of agents’ ability to process, reason about, and act on streams of information.
-Step RL for Creative Reasoning.
-Introducing reinforcement learning into the EscapeBench task is expected to enhance the accuracy and efficiency of the agent’s exploration in long-chain tasks.
-This implies that under the guidance of rewards and penalties, the agent can explore in the correct direction more quickly.
-Compared with existing end-to-end reinforcement learning schemes, which rely on the final completion of the escape room task as the ultimate reward, introducing step rewards—providing immediate feedback for each step of the model’s operation—could potentially accelerate convergence and foster creative advancements in the model.
-Specifically, the discrete steps in EscapeBench can be organized either as a continuous, logically coherent progression or as strategic, abrupt jumps reflecting non-linear reasoning.
-Step Rewards can also further draw on the idea of task decomposition to be structured hierarchically.
-By constructing such a framework, Hierarchical Reinforcement Learning could manage super-long reasoning chains, decompose complex tasks into manageable subtasks, and enable a more fine-grained exploration and evaluation of AI’s creative capabilities.
+Human-AI collaboration in EscapeBench may promote a new problem-solving paradigm by merging human intuition with AI’s systematic reasoning.
+Humans bring unique insights and ideas that AI might not generate, while AI excels in tasks like information aggregation and logical organization.
+This synergy fosters innovative strategies, improves efficiency, and creates opportunities for deeper learning, offering a dynamic and enriched problem-solving experience that bridges human creativity with AI’s structured problem-solving capabilities.
 8 Conclusion
 In this work, we introduce EscapeBench, the first benchmark for advancing LM’s creativity.
-Our results highlight that current LMs still lag in creative reasoning.
-However, the EscapeAgent framework effectively lowers the barrier to creative problem-solving and implicit goal identification, resulting in significant performance improvements.
-While these advancements are promising, challenges remain in further enhancing the creative capabilities of the language model itself.
-Future work may focus on integrating multi-modal perception and developing new reinforcement learning algorithms to inspire even greater creativity.
+Our results show that while LMs still lag in creative reasoning, the EscapeAgent framework improves innovative problem-solving and implicit goal identification.
+Despite these advancements, enhancing the models’ intrinsic creativity remains a challenge.
+Future work could explore integrating multi-modal perception and new reinforcement learning algorithms to foster greater creativity.
 Our work serves as an important first step, offering a robust environment for experimentation.
@@
 We anticipate that more powerful vision-language models may enable better automatic annotation in the future, though current model capabilities are still a limiting factor.
+Ethical Statement
+In this research, we consider the following ethical issues related to our benchmark and agent design:
+$\bullet$ Fairness: We ensure that EscapeBench is designed to provide equal evaluation opportunities for all agents, regardless of their underlying model architectures or training methodologies.
+The tasks and scenarios are crafted to assess creativity and problem-solving abilities without bias, promoting fairness in the benchmarking process.
+Additionally, we aim to avoid overfitting to specific agent strategies, ensuring a more generalizable and inclusive evaluation framework for future AI advancements.
+While our environment is robust, we caution against potential misuse and strongly encourage its fair and responsible use.
+$\bullet$ Transparency: Our work incorporates Chain-of-Thought reasoning in the BaseAgent framework to improve the transparency and interpretability of the agent’s decision-making process.
+This approach makes it easier to attribute the reasoning behind each agent’s action.
+Additionally, we will fully release the benchmarking code, EscapeAgent design, and data to promote transparency in our evaluation process, ensuring that the broader research community can benefit from and build upon our work.
+Acknowledgment
+This research is based upon work supported by U.S.
+DARPA ITM Program No.
+FA8650-23-C-7316 and DARPA ECOLE Program No.
+#HR00112390060.
+The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of DARPA, or the U.S.
+Government.
+The U.S.
+Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
 Appendix
@@
 The game engine involves scenes, tools, and items as three main components. We will introduce in detail each of them in the following.
+Scene.
 A scene typically includes a description, its connections to other scenes, and the tools and items it contains. A typical scene example in the game configuration looks like the following:
@@
 It leads to nearby scenes including “blocked path close-up” and “cabinet close-up”, where the model could reach through action “move(To the blocked path close-up)” and “move(To the cabinet close-up)”.
+Tool.
 Each tool has various states and a visibility status.
@@
 In this example, a rusty key (tool, state 1) hidden in the chest (item) won’t be visible to the user until the chest is opened, and after applying lubricant (tool), it may change to a non-rusty functional key (tool, state 2) that could be applied to open a safe.
+Item.
 Item is an upgraded version of the tool, as each state may await multiple inputs or tools in order to trigger certain changes, including item and tool’s visibility, state, etc.
@@
 These metrics vary significantly depending on the specific open-source models being tested.
-D.3 More Study Results
+Figure 12: Ablation of difficulty through Hints Used.
+Figure 13: Ablation of difficulty through Total Steps.
+D.3 Ablation Study
+We perform an ablation study on the total steps and used hints in Figure 12 and Figure 13.
+Generally, harder game settings require more steps and hints for an agent to solve.
+Since our difficulty setting depends solely on the granularity and usefulness of descriptions and feedback (see Table 1), our results demonstrate that the way the environment is presented can impact difficulty, even when the core game logic remains unchanged.
+Figure 14: A comparison of action trial times distribution for a specific item before success.
+D.4 Further Analysis
+We further analyze the valid actions agents attempt on different items in scenes before successfully operating them in Figure 14.
+While the BaseAgent exhibits a higher first-attempt success rate, EscapeAgent achieves greater effectiveness by making multiple attempts, leading to a significantly higher overall success rate within 10 trials.
+This difference can be attributed to EscapeAgent’s strategy of proposing multiple viable actions simultaneously.
+Although this increases the likelihood of eventually succeeding by trying more actions, it does not prioritize the most-likely-to-succeed action first.
+As a result, the BaseAgent appears more efficient on its initial trial, but its superficial approach, as highlighted in Table 4, limits its overall performance.
+The EscapeAgent’s design effectively addresses this limitation by leveraging a more exploratory approach, which proves advantageous in complex scenarios requiring creativity.
+D.5 More Study Results
 In Figure 9, we present additional results on the action step intervals with respect to two types of progress-making ways, achieving Key Step or Tool Collection.
@@
 Furthermore, for smaller models like Phi-3 and Ministral, EscapeAgent occasionally outperforms BaseAgent even in terms of one-trial success rates.
-This highlights how our framework effectively lowers the barriers for creative reasoning, even for less capable language models.
+This highlights how our framework effectively lowers the barriers to creative reasoning, even for less capable language models.
 In Figure 15, we showcase eight additional pairs of progress-making maps for eight more models.
@@
 Figure 15: More model’s analysis on progress with respect to action steps, extension of Figure 7.
+Appendix E Human Performance Details
+To better understand the “Average Human” baseline we present in the main table, we conducted a post-experiment survey with all participants, including:
+  • Prior offline Room Escape game experience? [Yes/No]
+  • Number of times played offline? [Integer]
+  • Prior online Room Escape experience (e.g., phone/PC)? [Yes/No]
+  • Self-assessed skill level? [Skilled / Might Be / Not Really / Not at All]
+  • Suggestions for improving EscapeBench
+Results indicated that nearly all participants had prior experience, with an average of 5.6 offline sessions.
+However, only 20% identified as skilled players; 40% selected “Might Be” and another 40% “Not Really.” Despite varying self-assessments, game logs showed no significant performance difference across groups—likely due to the trial-and-error nature of the tasks.
+Although a large-scale user study is beyond our current scope, these findings suggest that our reported human performance offers a fair approximation of average human reasoning under novel conditions.
+Appendix F Scalability of Benchmark Creation
+To ensure high-quality annotations, we employ human annotators supported by an automated checker and manual cross-validation.
+While we experimented with GPT-4o for annotation, it faced two major issues: (i) frequent omission of interactable objects, and (ii) inability to balance difficulty across scenarios.
+These limitations currently require human intervention to maintain benchmark quality.
+Specifically, we explored using GPT-4o to generate item-based scenarios.
+While the model can produce coherent logical sequences, it lacks the creativity and complexity required by our benchmark.
+For example:
+$\bullet$ Example 1: Use a magnifying glass to examine an ancient scroll, revealing hidden formulas.
+$\bullet$ Example 2: Clean a rusty key with water, then scrape it with a coin to unlock a cabinet (partially creative).
+Each AI-generated scenario yields about 12 key steps, whereas our benchmark includes over 1200 steps across 36 finely tiered scenarios.
+Therefore, human expertise remains essential for crafting creatively challenging and balanced content.
+While full automation falls short, parts of the process (e.g., validation, formatting) can be scaled.
+Future model improvements may increase automation potential, but human oversight will remain critical.
+Appendix G More Detailed Discussions
+We present a more detailed version of our results discussions and future research directions.
+Theoretical Foundations for AI Creativity.
+Understanding the cognitive mechanisms behind human creativity is essential for designing AI systems that emulate or surpass human creative processes.
+Human creativity is a multifaceted phenomenon involving the generation of novel and valuable ideas, problem-solving, and adaptation to complex situations Dainys (2024).
+Neurologically, it arises from the interplay between stochastic neuronal noise and structured, learned information, driven by spontaneous brain activity fluctuations Malach (2024).
+This contrasts with AI’s reliance on data patterns and algorithmic processes.
+Boden identifies three AI creativity mechanisms: combining familiar ideas in novel ways, exploring conceptual spaces, and enabling transformative innovations Boden (1998).
+Our research shows that even advanced models like GPT-4o struggle with implicit goal identification and creative problem-solving, often requiring extensive prompting.
+The EscapeAgent framework significantly reduces this dependency and improves task-solving efficiency, indicating that these modules effectively overcome barriers to creative reasoning.
+However, the results also highlight the importance of the core model’s capabilities, as larger models like GPT-4o benefit more from the framework than smaller models.
+This suggests further research is needed to address current models’ limitations in generating truly novel ideas.
+Interdisciplinary approaches integrating psychology, neuroscience, and model architecture can advance agent creativity further.
+Multimodal Integration.
+Expanding the escape room environment to include multimodal data, such as visual and voice cues, offers a promising avenue for enhancing both the agent’s performance and the realism of the scenarios.
+While integrating vision-language models could enable agents to interpret visual clues more naturally, such an extension would also require robust visual understanding and reasoning capabilities to handle the complexity of tasks effectively.
+Additionally, incorporating multimodal interactions presents opportunities to study how agents synthesize information across modalities, such as correlating visual patterns with instructions or adapting strategies based on dynamic, multimodal feedback.
+Future work should explore how to seamlessly integrate and harmonize these diverse data types into the benchmark, pushing the boundaries of agents’ ability to process, reason about, and act on streams of information.
+Step RL for Creative Reasoning.
+Introducing reinforcement learning into the EscapeBench task is expected to enhance the accuracy and efficiency of the agent’s exploration in long-chain tasks.
+This implies that under the guidance of rewards and penalties, the agent can explore in the correct direction more quickly.
+Compared with existing end-to-end reinforcement learning schemes, which rely on the final completion of the escape room task as the ultimate reward, introducing step rewards—providing immediate feedback for each step of the model’s operation—could potentially accelerate convergence and foster creative advancements in the model.
+Specifically, the discrete steps in EscapeBench can be organized either as a continuous, logically coherent progression or as strategic, abrupt jumps reflecting non-linear reasoning.
+Step Rewards can also further draw on the idea of task decomposition to be structured hierarchically.
+By constructing such a framework, Hierarchical Reinforcement Learning could manage super-long reasoning chains, decompose complex tasks into manageable subtasks, and enable a more fine-grained exploration and evaluation of AI’s creative capabilities.